### v1

# A Comprehensive Survey on the Trustworthiness of Large Language Models in Healthcare

Abstract
The application of large language models (LLMs) in healthcare has the potential to revolutionize clinical decision-making, medical research, and patient care. As LLMs are increasingly integrated into healthcare systems, several critical challenges must be addressed to ensure their reliable and ethical deployment. These challenges include truthfulness, where models generate misleading information; privacy, with risks of unintentional data retention; robustness, requiring defenses against adversarial attacks; fairness, addressing biases in clinical outcomes; explainability, ensuring transparent decision-making; and safety, mitigating risks of misinformation and medical errors. Recently, researchers have begun developing benchmarks and evaluation frameworks to systematically assess the trustworthiness of LLMs. However, the trustworthiness of LLMs in healthcare remains underexplored, lacking a systematic review that provides a comprehensive understanding and future insights into this area. This survey bridges this gap by providing a comprehensive overview of the recent research of existing methodologies and solutions aimed at mitigating the above risks in healthcare. By focusing on key trustworthiness dimensions including truthfulness, privacy and safety, robustness, fairness and bias, and explainability, we present a thorough analysis of how these issues impact the reliability and ethical use of LLMs in healthcare. This paper highlights ongoing efforts and offers insights into future research directions to ensure the safe and trustworthy deployment of LLMs in healthcare.
A Comprehensive Survey on the Trustworthiness of Large Language Models in Healthcare
Manar Aljohani1, Jun Hou1, Sindhura Kommu1, Xuan Wang1, 1Department of Computer Science, Virginia Tech, Blacksburg, VA, USA (manara, junh, sindhura, xuanw)@vt.edu
1 Introduction
The application of LLMs in healthcare is advancing rapidly, with the potential to transform clinical decision-making, medical research, and patient care. However, incorporating them into healthcare systems poses several key challenges that need to be addressed to ensure their reliable and ethical use. As highlighted in Bi et al. (2024), a major concern is the trustworthiness of AI-enhanced biomedical insights. This encompasses improving model explainability and interpretability, enhancing robustness against adversarial attacks, mitigating biases across diverse populations, and ensuring strong data privacy protections. Key concerns include truthfulness, privacy, safety, robustness, fairness, and explainability, each of which plays a vital role in the reliability and trustworthiness of AI-driven healthcare solutions.
Truthfulness, defined as "the accurate representation of information, facts, and results by an AI system" Huang et al. (2024), is critical in healthcare, as inaccuracies can lead to misdiagnoses or inappropriate treatment recommendations. Ensuring that generated information is both accurate and aligned with verified medical knowledge is essential. Additionally, privacy concerns arise from the risk of exposing sensitive patient data during model training and usage, potentially leading to breaches or violations of regulations such as HIPAA (Health Insurance Portability and Accountability Act) and GDPR (General Data Protection Regulation). Ensuring patient confidentiality while leveraging LLMs for diagnostics and treatment recommendations is a critical challenge. Safety, defined as “ensuring that LLMs do not answer questions that can harm patients or healthcare providers in healthcare settings” Han et al. (2024b), further underscores the necessity of implementing stringent safeguards to mitigate harm. Robustness refers to an LLM’s ability to consistently generate accurate, reliable, and unbiased outputs across diverse clinical scenarios while minimizing errors, hallucinations, and biases. It also encompasses the model’s resilience against adversarial attacks, ensuring that external manipulations do not compromise its integrity. A truly robust LLM in healthcare must demonstrate stability, reliability, and fairness, even when faced with noisy, ambiguous, or adversarial inputs, thereby safeguarding patient safety and supporting clinical decision-making. Similarly, fairness and bias must be addressed to prevent discriminatory patterns in model predictions, which could lead to unequal treatment recommendations and exacerbate healthcare disparities. Furthermore, the explainability of LLMs, which ensures that model outputs are interpretable and transparent, plays a vital role in fostering trust and allowing informed decision-making by healthcare professionals. The lack of transparency in model reasoning complicates clinical adoption and raises concerns about accountability.
Tackling these challenges is essential for the trustworthy and ethical implementation of LLMs in healthcare. Recently, researchers have begun developing benchmarks and evaluation frameworks to systematically assess the trustworthiness of LLMs Huang et al. (2024). The trustworthiness of LLMs in healthcare is gaining increasing attention due to its significant social impact. However, there is currently no systematic review that provides a comprehensive understanding and future insights into this area. To bridge this gap, we present a comprehensive survey that explores these trust-related dimensions in detail, reviewing existing datasets, solutions, and methodologies aimed at improving the trustworthiness of LLMs in healthcare.
2 Datasets, Models, and Tasks
We first conducted an extensive search for papers on the trustworthiness of LLMs in healthcare. Our search utilized a range of keyword combinations, including terms such as ‘large language models,’ ‘foundation model,’ ‘medical,’ ‘clinical,’ ‘explainability,’ ‘truthfulness,’ ‘trustworthiness,’ ‘safety,’ ‘fairness,’ ‘robustness,’ and ‘privacy.’ We explored several reputable venues, including Arxiv, PubMed, ACL, EMNLP, NAACL, ICML, NeurIPS, ICLR, KDD, Nature, Science, AAAI, and IJCAI, with a focus on recent publications post-2021. After reviewing the search results, we identified a total of 30,595 papers. Following the removal of duplicates, we narrowed the focus to 69 papers that specifically addressed the truthfulness, privacy, safety, robustness, fairness, bias, and explainability of LLMs in the healthcare domain.
We then summarized all the datasets, models, and tasks relevant to research on trust in LLMs for healthcare, providing a comprehensive overview of their applications and contributions to this domain. The datasets used in studies of trust in LLMs for healthcare are categorized by the dimensions of trustworthiness they address in Appendix A, where we highlight key details such as data type, content, task, and dimensions of trustworthiness. The content of each dataset specifies its composition, while the task refers to the primary purpose for which the dataset is utilized. The data type varies across datasets and includes web-scraped data, curated domain-specific datasets, public text corpora, synthetic data, real-world data, and private datasets, providing a comprehensive overview of their relevance to healthcare applications. The models assessed in studies on trust in LLMs for the healthcare domain are outlined, along with their trustworthiness dimensions, in Appendix B, where we summarized key details such as the model name, release year, task, and the institution responsible for its development. The tasks covered various primary focuses of LLMs in healthcare. Based on insights from the survey by Liu et al. (2024), these tasks are outlined as follows:
Medical Information Extraction (Med-IE)
Med-IE extracts structured medical data from unstructured sources such as EHRs, clinical notes, and research articles. Key tasks include entity recognition (identifying diseases, symptoms, and treatments), relationship extraction (understanding entity connections), event extraction (detecting clinical events and attributes), information summarization (condensing medical records), and adverse drug event detection (identifying medication-related risks).
Medical Question Answering (Med-QA)
Med-QA systems interpret and respond to complex medical queries from patients, clinicians, and researchers. Their core functions include query understanding (interpreting user questions), information retrieval (finding relevant data in medical databases), and inference and reasoning (drawing conclusions, inferring relationships, and predicting outcomes based on retrieved data).
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=14em,font=,, where level=2text width=18em,font=,, where level=3text width=18em,font=,, [ Trustworthiness of LLMs in Healthcare, ver [ Truthfulness, fill=blue!10 [ Med-HALT Pal et al. (2023), Med-HVL Yan et al. (2024b), Med-HallMark Chen et al. (2024), Semantic Entropy Farquhar et al. (2024),
SEPs Han et al. (2024a), Survey Ahmad et al. (2023), Self Reflection Ji et al. (2023), Model-Agnostic Hallucination Post-Processing Li et al. (2024),
Faithful Reasoning Tan et al. (2024), PubHealthTab Akhtar et al. (2022), HEALTHVER Sarrouti et al. (2021), CRITIC Gou et al. (2024),
Cross-Examination Cohen et al. (2023), KnowledgeEditor De Cao et al. (2021), leaf, text width=66em ] ] [ Privacy, fill=blue!10 [ Survey Das et al. (2024), Survey Yan et al. (2024a), Privacy Risks Pan et al. (2020), Federated Learning Zhao et al. (2024),
Differential Privacy Singh et al. (2024), SecureSQL Song et al. (2024), Memorize Fine-tuning Data Yang et al. (2024a),
clinical Note De-identification Altalla’ et al. (2025), De-identification Liu et al. (2023b), leaf, text width=66em ] ] [ Safety, fill=blue!10 [ Med-harm Han et al. (2024c), Medsafetybench Han et al. (2024b), UNIWIZ Das and Srihari (2024), Misinformation Attacks Han et al. (2024d),
MEDIC Kanithi et al. (2024), leaf, text width=66em ] ] [ Robustness, fill=blue!10 [ Survey Yuan et al. (2023), Survey Alberts et al. (2023), Stumbling Blocks Wang et al. (2024), LLM-TTA O’Brien et al. (2024), MedFuzz Ness et al. (2024),
Detecting Anomalies Rahman et al. (2024), Instruction Phrasings Ceballos-Arroyo et al. (2024), Adversarial Attacks Yang et al. (2024b),
Secure Your Model Tang et al. (2024), leaf, text width=66em ] ] [ Fairness and Bias, fill=blue!10 [ Evaluation StudyZack et al. (2024), BiasMedQA Schmidgall et al. (2024), Survey Parray et al. (2023), Instruction Fine-tuning Singhal et al. (2023),
Hurtful Words Zhang et al. (2020), Race-based Medicine Omiye et al. (2023), Detect Debunked Stereotypes Swaminathan et al. (2024),
EquityMedQA Pfohl et al. (2024), Superficial Fairness Alignment Wei et al. (2024), Examines Biased AI Adam et al. (2022),
Identify Biases Yang et al. (2024c), Quantifying Cognitive Biases Lin and Ng (2023), Mitigate Cognitive Biases Ke et al. (2024), CI4MRC Zhu et al. (2023), leaf, text width=66em ] ] [ Explanability, fill=blue!10 [ Knowledge Graphs Shariatmadari et al. (2024), LLMs and Explainable ML Elsborg and Salvatore (2023), Medical Imaging Explainability Ghosh et al. (2023),
MedViLaM Xu et al. (2024), MedThink Gai et al. (2024), MedExQA Kim et al. (2024), Causal Graphs Meet Thoughts Luo et al. (2025),
Retrieval and Reasoning on KGs Ji et al. (2024), TOSRR Liu et al. (2025), DDCoT Zheng et al. (2023), Layered Chain-of-Thought Prompting Sanwal (2025),
A ChatGPT Aided Explainable Framework Liu et al. (2023a), leaf, text width=66em ] ] ]
Figure 1: Summary of the recent research across various dimensions of trustworthiness of LLMs in healthcare.
Medical Natural Language Inference (Med-NLI)
Med-NLI analyzes the logical relationships between medical texts. Key tasks include textual entailment (determining if one statement logically follows another), contradiction detection (identifying conflicting statements), neutral relationship identification (recognizing unrelated statements), and causality recognition (inferring cause-and-effect relationships).
Medical Text Generation (Med-Gen)
Med-Gen focuses on generating and summarizing medical content. Its key applications include text summarization (condensing lengthy documents into concise summaries) and content generation (producing new medical descriptions or knowledge based on input data).
3 Trustworthiness of LLMs in Healthcare
We examine the challenges related to the trustworthiness of LLMs in healthcare, outlining key strategies for identifying and mitigating these concerns. From our literature review screening, we identified truthfulness, privacy and safety, robustness, fairness and bias, and explainability as key trustworthiness dimensions of LLMs as highlighted in TrustLLM Huang et al. (2024), particularly in healthcare. Figure 1 provides a summary of the recent research on trust in LLMs for healthcare across key dimensions of trustworthiness.
3.1 Truthfulness
Ensuring the truthfulness of LLMs in healthcare is vital, as inaccurate information can significantly impact patient care and clinical outcomes. Given their influence on diagnoses and treatment decisions, it is essential to develop effective methods to detect and mitigate hallucinations and factual inaccuracies.
Hallucinations in medical LLMs arise from reliance on unverified sources, biases in training data, and limitations in contextual understanding and sequential reasoning Ahmad et al. (2023). Addressing these issues requires robust evaluation frameworks, self-correction mechanisms, and uncertainty quantification techniques.
The Med-HALT benchmark Pal et al. (2023) is designed to evaluate hallucinations in medical LLMs by using reasoning-based tests like ‘False Confidence’ and ‘None of the Above,’ as well as memory-based tests to assess how well the model recalls medical knowledge. On the other hand, the interactive self-reflection methodology Ji et al. (2023) aims to reduce hallucinations in medical question-answering tasks by introducing an iterative feedback loop where the model refines its responses through self-evaluation and knowledge adjustment.
In the context of multimodal models, Med-HVL Yan et al. (2024b) introduces two key metrics—Object Hallucination and Domain Knowledge Hallucination—to quantify hallucinations in Large Vision-Language Models (LVLMs). This framework also uses the CHAIR (Caption Hallucination Assessment with Image Relevance) metric to assess object hallucinations in image captioning. Med-HallMark Chen et al. (2024) takes it a step further by providing a multi-task evaluation framework with a hierarchical categorization of hallucinations, introducing the MediHall Score for assessing hallucination severity and the MediHallDetector, a multitask-trained LVLM for hallucination detection.
Researchers have also investigated semantic entropy, a probabilistic measure of uncertainty, to detect hallucinations in LLMs. For example, Farquhar et al. (2024) leverage semantic entropy to identify confabulations—hallucinations where the model generates arbitrary or incorrect outputs. While effective, this approach is computationally expensive, limiting its scalability. To overcome this, Han et al. (2024a) introduce Semantic Entropy Probes (SEPs), which approximate semantic entropy directly from hidden states. By eliminating the need for multiple output samples, SEPs significantly reduce computational overhead. Both methods have been successfully applied to biomedical datasets, such as BioASQ.
Although these techniques offer valuable contributions, hallucination mitigation methods often lack adaptability, being either task-specific or requiring expensive retraining. To address this gap, MEDAL Li et al. (2024) introduces a model-agnostic post-processing framework that integrates with any medical summarization model. MEDAL uses a self-examining correction model to improve factual accuracy without adding extra computational costs, providing a practical solution to the issue of medical hallucinations.
Collectively, these studies underscore the multifaceted challenge of ensuring truthfulness in medical LLMs. By leveraging benchmarking frameworks, self-correction mechanisms, and entropy-based uncertainty measures, researchers can develop complementary strategies for detecting, mitigating, and quantifying hallucinations. A key focus of these efforts is the development of quantifiable scoring methods, enabling systematic assessment and comparison across different models. These evaluation techniques not only help identify the most reliable and effective LLMs for healthcare applications but also provide actionable insights for further improvements.
Factual accuracy is fundamental to building trust in LLMs, especially in healthcare, where reliable and verifiable information is critical. However, current LLMs lack effective mechanisms to trace claims back to their original sources, underscoring the urgent need for improved validation techniques to ensure safe and trustworthy medical applications. To address these challenges, several studies have introduced innovative approaches to enhance the transparency, accuracy, and reliability of healthcare LLMs.
Tan et al. (2024) propose an approach that integrates multiple perspectives from scientific literature to evaluate conflicting arguments, thereby improving LLM reasoning. Similarly, Akhtar et al. (2022) introduce PubHealthTab, a table-based dataset designed for validating public health claims against noisy evidence, while Sarrouti et al. (2021) present HEALTHVER, a dataset tailored for evidence-based fact-checking of health-related claims. These structured benchmarks provide a foundation for assessing and refining the reliability of LLM-generated medical information.
Beyond dataset-driven validation, self-correction mechanisms have been explored to improve LLM truthfulness. Gou et al. (2024) introduce CRITIC, a framework inspired by human fact-checking practices, enabling LLMs to validate and refine their responses through iterative feedback and evaluation. Expanding on automated fact-checking, Cohen et al. (2023) propose a cross-examination framework, where an examiner LLM identifies inconsistencies through multi-turn interactions with the original model. Unlike fully automated verification pipelines, CRITIC incorporates human-like evaluation strategies, enhancing the trustworthiness of fact-checking in medical contexts.
Overall, these studies advance the factual accuracy and transparency of LLMs in healthcare by introducing structured benchmarks, iterative validation processes, and automated fact-checking strategies. By incorporating these approaches, researchers can enhance the reliability of medical LLMs, ensuring they deliver more accurate, evidence-based insights to support clinical decision-making.
3.2 Privacy
LLM-based healthcare applications pose significant privacy risks due to their ability to memorize and reproduce sensitive patient data Das et al. (2024). Unauthorized data exposure can lead to confidentiality breaches, ethical concerns, and compliance violations Pan et al. (2020). Addressing these risks requires privacy safeguards at different stages of model development.
A major challenge is unintended data retention and leakage, where LLMs memorize fine-tuning data, increasing re-identification risks. Studies show that domain-specific LLMs, such as Medalpaca, can retain sensitive data, making privacy breaches more likely Yang et al. (2024a). Additionally, adversarial attacks like prompt injection and inference attacks can further exploit these vulnerabilities, as demonstrated by the SecureSQL benchmark Song et al. (2024).
To mitigate these risks, pre-training privacy safeguards focus on de-identification. Altalla’ et al. (2025) assess GPT-3.5 and GPT-4 in clinical note de-identification and synthetic data generation. Similarly, Liu et al. (2023b) propose a GPT-4-enabled framework for masking private information while maintaining text structure. However, de-identification remains imperfect, as attackers may infer sensitive details from anonymized text.
During fine-tuning, techniques such as federated learning Zhao et al. (2024) and differential privacy Singh et al. (2024), as highlighted by Liu et al. (2024), play a crucial role in safeguarding patient data. Federated learning enables decentralized training without sharing raw data, but it demands high computational resources. Differential privacy adds noise to protect sensitive information but can reduce model accuracy.
Adversarial defenses remain limited. The SecureSQL benchmark Song et al. (2024) highlights LLM vulnerabilities to structured query attacks. While chain-of-thought (COT) prompting offers partial mitigation, it does not eliminate the risk of data exposure.
Researchers address privacy concerns in healthcare LLMs through two primary approaches: de-identification, which alters real data to prevent re-identification, and synthetic data generation, which creates artificial data to eliminate reliance on sensitive patient information. While these strategies enhance privacy protection and maintain model effectiveness, challenges remain in long-term memorization control and adversarial robustness, requiring further research to strengthen data security and prevent unintended information retention.
3.3 Safety
Ensuring the safety of LLMs in healthcare is critical, as these models must not generate harmful responses. A key safety concern, noted by Han et al. (2024d), is that modifying just 1.1 % of a model’s weights can embed persistent biomedical inaccuracies while maintaining overall performance. This highlights the need for rigorous validation mechanisms and safety assessments to prevent misleading medical information before clinical use.
To systematically assess safety, MedSafetyBench Han et al. (2024b) was introduced as the first benchmark designed to evaluate LLM safety in medical contexts. It includes 1,800 harmful medical queries alongside safety-optimized responses generated using advanced LLMs and adversarial techniques. Results indicate that publicly available medical LLMs fail to meet safety standards, but fine-tuning with MedSafetyBench significantly improves safety without compromising performance.
A major challenge is that adversarial actors can manipulate LLMs to generate unsafe outputs, while excessive safety alignment may induce hallucinations. To address this, Das and Srihari (2024) propose UNIWIZ, a two-step framework that unifies safety alignment and factual knowledge retrieval. Their safety-priming approach synthesizes safety-focused training data, while a retrieval mechanism ensures that model outputs remain factually accurate. Models fine-tuned on UNIWIZ outperform larger state-of-the-art instruction-tuned models across multiple safety and accuracy metrics.
Another key contribution to safety alignment is from Han et al. (2024c), who provide the first comprehensive safety evaluation for medical LLMs (MedLLMs). They define key concepts of medical safety and alignment and introduce Med-Harm, a dataset designed to evaluate both general and medical-specific risks. This dataset assesses how well LLMs handle harmful medical questions, ensuring they adhere to safety and ethical standards in medical AI.
Further advancing safety assessments, Kanithi et al. (2024) introduce MEDIC, a multi-dimensional trustworthiness evaluation framework that systematically assesses medical LLMs across critical dimensions of clinical competence including medical reasoning and clinical safety.
These studies collectively offer a multi-faceted approach to LLM safety in healthcare. MedSafetyBench Han et al. (2024b) provides a standardized benchmark for safety evaluation and fine-tuning, while UNIWIZ Das and Srihari (2024) introduces a structured framework that prevents hallucinations while reinforcing safety. Han et al. (2024c) focus on comprehensive safety alignment, establishing Med-Harm to evaluate domain-specific risks. Lastly, MEDIC Kanithi et al. (2024) offers a holistic evaluation franmework for improving practical application of LLMs in clinical settings. Together, these efforts contribute to a more rigorous and systematic framework for assessing and improving LLM safety in medical applications.
3.4 Robustness
Enhancing the robustness of LLMs is crucial for their reliability in healthcare applications. A key approach involves developing adversarial test samples tailored to the medical domain, such as synthetic anomaly cases Yuan et al. (2023) and boundary stress testing Wang et al. (2024), to assess model resilience. However, creating clinically meaningful adversarial samples presents unique challenges, as Alberts et al. (2023) highlight the need to align adversarial testing methods with the complexities of real-world medical data, where medical dependencies must be accounted for.
In addition to adversarial testing, uncertainty quantification is another important avenue to improve robustness. LLM-TTA O’Brien et al. (2024) explores test-time adaptation techniques to enhance model performance on rare or unfamiliar cases, which are common in medical diagnostics. Unlike adversarial robustness, which focuses on resistance to manipulated inputs, uncertainty quantification aims to identify when models are likely to be incorrect, providing a complementary safety mechanism.
Another critical question is whether benchmark performance truly reflects medical robustness. MedFuzz Ness et al. (2024) challenges assumptions in MedQA by modifying questions to test if models rely on rigid, dataset-specific patterns rather than genuine clinical reasoning. This research exposes vulnerabilities in LLMs, revealing that subtle changes in input can significantly impact performance, raising concerns about their reliability in dynamic medical environments.
Instruction robustness is also a growing concern. Ceballos-Arroyo et al. (2024) examine how variations in medical instructions affect performance across different LLMs, finding that specialized medical models may be more fragile than general-purpose models when instructions are reworded. This counterintuitive result suggests that excessive domain adaptation may decrease flexibility and reduce robustness.
Adversarial vulnerabilities also pose direct security risks. Yang et al. (2024b) investigate two adversarial attack strategies across medical tasks using real patient data, demonstrating that fine-tuned models are especially vulnerable to poisoning attacks that subtly alter learned weights. While adversarial data does not always degrade general performance, it can introduce dangerous biases into specific medical predictions, making early attack detection a priority.
To protect against adversarial manipulations, Tang et al. (2024) introduce Secure Your Model, a framework that strengthens LLM robustness with cryptographic prompt authentication. This mechanism ensures that only verified and secure prompts are processed, mitigating vulnerabilities associated with prompt injections and adversarial attacks, and reducing the risk of model exploitation in medical contexts.
While all these studies address LLM robustness, they differ in their primary focus. MedFuzz Ness et al. (2024) and Ceballos-Arroyo et al. (2024) expose vulnerabilities in existing benchmarks and task instructions, questioning whether current evaluation methods truly measure robustness or merely reflect dataset biases. In contrast, Yang et al. (2024b) and Alberts et al. (2023) highlight adversarial threats, demonstrating how medical LLMs can be subtly manipulated, raising concerns about their security in real-world applications. Meanwhile, LLM-TTA O’Brien et al. (2024) takes a different approach, focusing on uncertainty quantification rather than adversarial resistance to enhance reliability in handling unfamiliar cases. Secure Your Model Tang et al. (2024) provides an additional security layer by introducing proactive adversarial defenses through prompt protection mechanisms, ensuring resilience against manipulation risks.
These studies highlight that robustness is a multifaceted challenge, requiring advancements in evaluation methods and defensive mechanisms. Ensuring LLMs can handle adversarial scenarios, integrate domain knowledge, and adapt to language variations is crucial for their safe deployment in healthcare. Strengthening robustness through testing and resilience-building enhances the trustworthiness of medical LLMs, making them more reliable in complex clinical settings.
3.5 Fairness and Bias
Ensuring fairness in LLMs is crucial in healthcare, where biased models can result in unequal treatment outcomes. Research has highlighted biases in clinical data and practice related to race, gender, and disability. For example, Omiye et al. (2023) examine the potential for harmful or inaccurate race-based content in LLMs, while Zack et al. (2024) discusse how language models encode societal biases that can affect healthcare outcomes. These studies underscore the need for fairness in LLM development to ensure equitable healthcare delivery.
Efforts to address bias focus on both detection and mitigation. Swaminathan et al. (2024) offers an automated method for detecting race-based medicine stereotypes, while BiasMedQA Schmidgall et al. (2024) benchmarks cognitive biases in medical tasks across multiple models, revealing varying bias resilience. Mitigation strategies, such as bias education, and one-shot and few-shot bias demonstrations, are proposed to reduce but not fully eliminate bias. Pfohl et al. (2024) introduce frameworks for assessing health equity-related harms in LLMs, including EquityMedQA, a dataset for equity-focused testing. Additionally, Wei et al. (2024) distinguishe between intrinsic fairness, rooted in model training, and behavioral fairness, which relates to model operation in real-world applications, advocating for both to ensure equitable outcomes.
Bias is also explored in closed-source models. Zack et al. (2024) evaluate racial and gender biases in clinical scenarios, finding that models often amplify societal biases. Similarly, Adam et al. (2022) show that biased AI recommendations can affect emergency decisions, while Yang et al. (2024c) identify healthcare disparities in model predictions based on patient demographics. For open-source LLMs, techniques such as reinforcement learning with clinician feedback Zack et al. (2024) and data augmentation during pre-training Parray et al. (2023) enhance training data quality to reduce bias. In contrast, for closed-source models, where internal representations are inaccessible, strategies like instruction fine-tuning Singhal et al. (2023) and prompt engineering Schmidgall et al. (2024) are employed to improve fairness in outputs.
Further, Zhang et al. (2020) investigate how LLM embeddings can encode biases, particularly in clinical tasks, and applies adversarial debiasing to mitigate disparities. Lin and Ng (2023) identify cognitive biases in BERT, while Ke et al. (2024) use a multi-agent framework to explore how LLMs can mitigate biases in clinical decision-making. Additionally, Zhu et al. (2023) introduce CI4MRC, a method that addresses name-related bias in Machine Reading Comprehension (MRC) tasks by applying a causal interventional paradigm.
In addressing bias, solutions vary depending on whether the bias is at the individual level (e.g., name-related information) or dataset level (e.g., biased racial distributions). Open-sourced LLMs benefit from direct accessibility, enabling robust interventions to mitigate bias. In contrast, closed-sourced models, due to their inaccessibility, rely on methods such as instruction fine-tuning and external post-processing tools to refine model outputs and reduce bias.
3.6 Explanability
The lack of explainability in LLMs poses a significant barrier to building trust with clinical practitioners, thereby restricting their adoption in real-world healthcare systems. To address this challenge, efforts in clinical practice and research have prioritized improving the transparency of LLMs by developing more effective explanation mechanisms and incorporating human oversight. For instance, Shariatmadari et al. (2024) enhance biomedical applications by integrating knowledge graphs with language models and visualizing attention probabilities to provide clear and interpretable explanations for model predictions. Elsborg and Salvatore (2023) employ local explanation models to generate intuitive, case-specific insights, further advancing the explainability of LLMs and fostering trust in their medical applications.
Another critical area of research focuses on medical imaging explainability. Ghosh et al. (2023) propose a method that iteratively decomposes a black-box (BB) model into interpretable expert models and a residual network, where expert models specialize in specific data subsets and explain their reasoning using First-Order Logic (FOL).
Efforts to improve reasoning and multimodal integration in LLMs have also gained momentum. MedViLaM Xu et al. (2024) and MedThink Gai et al. (2024) focus on improving medical understanding and reasoning by integrating both textual and visual data for complex medical tasks such as question answering and medical image classification, aiming to provide a more holistic view for decision-making. In contrast, MedExQA Kim et al. (2024) prioritizes explainability in medical QA systems by providing multiple explanations for its responses. Causal Graphs Meet Thoughts Luo et al. (2025) and Retrieval and Reasoning on KGs Ji et al. (2024) explore how to enhance complex reasoning in LLMs through knowledge graphs (KGs), helping these models retrieve relevant information from structured sources to answer questions more effectively and with better justification.
For more specialized applications, TOSRR Liu et al. (2025) introduce tree-organized self-reflective retrieval, which aims to improve performance in the niche area of Traditional Chinese Medicine (TCM), combining LLMs with self-reflection techniques for more accurate and contextually relevant responses. Studies like DDCoT Zheng et al. (2023) and Layered Chain-of-Thought Prompting Sanwal (2025) develop methods for multimodal reasoning across multiple agents, with DDCoT focusing on chain-of-thought prompting for structured reasoning across tasks. Lastly, frameworks like A ChatGPT Aided Explainable Framework Liu et al. (2023a) aim for zero-shot diagnosis and interpretation, ensuring more accessible AI support in real-world medical settings.
The trend towards explainability enhancement encompasses both intrinsic and post-hoc approaches. Intrinsic methods, such as knowledge graphs and decomposed expert models, integrate explainability directly into the models, offering natural transparency. Post-hoc methods, including local explanations and textual justifications, elucidate model decisions after predictions, improving user comprehension and trust.
4 Future Directions
We have reviewed key trust challenges in LLMs and existing solutions. This section highlights current limitations and proposes future directions.
While efforts to enhance LLM truthfulness in healthcare have advanced, gaps remain, including limited adaptability in hallucination mitigation and weak source attribution. Future research should prioritize model-agnostic post-processing, improved self-correction, enhanced uncertainty quantification, and real-time fact-checking via structured knowledge integration.
Existing privacy safeguards, such as de-identification and federated learning, remain imperfect. Strengthening de-identification methods, reinforcing federated learning defenses, refining differential privacy, and exploring homomorphic encryption and real-time audits are crucial next steps.
Safety remains a pressing concern due to adversarial attacks and excessive safety alignment-induced hallucinations. Future work should focus on robust validation mechanisms, refined safety alignment strategies, and comprehensive evaluation frameworks.
Improving robustness remains critical, with adversarial testing and uncertainty quantification methods needing to better handle medical data complexities. Future research should focus on clinically relevant adversarial tests, enhancing uncertainty techniques, and improving instruction robustness.
Despite progress in fairness, key gaps persist, including the need for comprehensive bias assessments and real-world testing of mitigation strategies. Standardizing fairness metrics, conducting real-world evaluations, and assessing long-term impacts on healthcare equity are critical for progress.
Advancements in explainability have yet to bridge significant gaps. Future research should focus on integrating explainability into clinical workflows, developing interactive explanations, and improving multimodal integration to enhance transparency and trust.
Recently, multi-agent frameworks like TriageAgent Lu et al. (2024) have been introduced to streamline complex clinical tasks through agent collaboration. By harnessing their capabilities, we can embed trustworthiness mitigation and evaluation within a multi-agent system, enabling proactive monitoring and intervention across key areas such as truthfulness, privacy, robustness, fairness, and explainability in healthcare LLMs.
Addressing these gaps will ensure that LLMs can be effectively integrated into healthcare systems, improving their reliability, privacy, safety, fairness, and transparency.
5 Conclusion
The integration of LLMs into healthcare holds great promise, but realizing their full potential requires addressing the critical challenges outlined in this survey. A key concern is truthfulness, as inaccuracies in medical LLMs pose serious risks to patient safety, making their detection and mitigation an ongoing research priority. Equally vital are privacy, safety, robustness, fairness, and explainability, ensuring responsible deployment in real-world clinical settings.
While existing solutions show progress, much work remains to enhance the reliability, transparency, and ethical implications of LLMs in healthcare. Future research must refine these areas, balancing performance with trustworthiness while preventing LLM-based systems systems from worsening healthcare disparities. Comprehensive benchmarks, cross-disciplinary collaboration, and model accountability frameworks will be essential. Additionally, regulatory oversight and ethical guidelines must ensure LLM applications align with medical standards and patient rights.
Ultimately, achieving safe and equitable AI-driven healthcare will require ongoing efforts to improve both technical capabilities and societal frameworks.
Limitations
This survey provides a comprehensive overview of the challenges associated with LLMs in healthcare, but it primarily focuses on existing methodologies, leaving out emerging technologies that could address these issues in new ways. It also lacks practical insights into the real-world implementation of these solutions, such as deployment challenges, cost considerations, and system integration, which would make the findings more applicable to healthcare settings.
While the paper addresses privacy and safety, it does not fully explore broader ethical issues like informed consent, patient autonomy, and human oversight. Additionally, the survey focuses on current research without delving into the long-term societal and health impacts of LLM deployment, such as changes in doctor-patient relationships, patient trust, and healthcare workflows.
Appendix A Comparison of Datasets
We systematically collected and analyzed 41 datasets relevant to the study of trust in LLMs for healthcare. Table 1 provides a comprehensive summary, highlighting key attributes such as data type, content, associated tasks, and the specific trustworthiness dimensions they address. These datasets vary widely, including web-scraped data, curated domain-specific datasets, public text corpora, synthetic data, real-world data, and private datasets. Each dataset’s content specifies its composition, while its associated task defines its primary research application. Additionally, we categorize the datasets based on critical trustworthiness dimensions—truthfulness, privacy and safety, robustness, fairness and bias, and explainability—offering a structured evaluation of their contributions to building reliable and trustworthy healthcare AI.
Datasets Data Type Content Task Dimensions
MultiMedQA Combination of Public and Synthetic Data, Curated Domain-Specific Dataset A benchmark combining six existing medical questions answering datasets spanning professional medicine, research and consumer queries and a new dataset of medical questions searched online, HealthSearchQA. Tasks including Medical Question Answering, Clinical Reasoning, Evidence-Based Medicine, Multilingual and Multimodal Support, Bias and Safety Analysis Fairness and Bias
BiasMedQA Curated Domain-Specific Datasets 1273 USMLE questions Replicate common clinically relevant cognitive biases Fairness and Bias
NEJM Healer Real Data and Curated Domain-Specific Dataset, Consists of Clinical Cases, Diagnostic Pathways, Educational Materials, Interactive Learning Modules Tasks including Diagnostic Skill Development, Medical Education, Simulated Decision-Making, Feedback and Improvement Fairness and Bias
EquityMedQA Curated domain-specific datasets and synthetic data Cover a wide range of medical topics to surface biases that could harm health equity, including implicit and explicit adversarial questions addressing biases like stereotypes, lack of structural explanations, and withholding information. Evaluate the performance of LLMs in generating unbiased, equitable medical responses. Fairness and Bias
”bias” medical dataset curated, domain-specific dataset comprising real-world healthcare data from a large U.S. health system Includes patient demographics, medical histories, and healthcare utilization records. Evaluate and identify racial biases in algorithms used for healthcare management. Fairness and Bias
SQuAD Curated Domain-Specific Dataset Consists of over 100,000 question-answer pairs derived from more than 500 articles from Wikipedia. Each question is paired with a segment of text from the corresponding article, serving as the answer. To develop models that can read a passage and answer questions about it, assessing the model’s ability to understand and extract information from the text. Fairness and Bias
MIMIC Real Data and Curated Domain-Specific Datasets Consists of electronic health records include patient Demographics, Clinical Data, Medical Notes, Treatment Records, Time-series Data Various medical and machine-learning tasks, including Clinical Decision Support, Disease Modeling, Natural Language Processing, Time-series Analysis, Education and Research Fairness and Bias, Explainability
MedQA Curated Domain-Specific Datasets A benchmark that includes questions drawn from the United States Medical License Exam (USMLE). Exam the physicians to test their ability to make clinical decisions Fairness and Bias, Robustness, Explainability
PMC-Patients Curated dataset derived from public text corpora. Contains 167,000 patient summaries extracted from 141,000 PMC articles Designed to benchmark ReCDS systems through two primary tasks: Patient-to-Article Retrieval (PAR), Patient-to-Patient Retrieval (PPR) Robustness
MIMIC- III Public text corpora, real-world data De-identified health-related data from over 40,000 critical care patients, including demographics, vital signs, laboratory tests, medications, and caregiver notes. Epidemiological studies, clinical decision-rule improvement, machine learning in healthcare. Robustness
Datasets Data Type Content Task Dimensions
MedSafetyBench Curated domain-specific dataset and synthetic (generated using GPT-4, Llama-2-7b-chat, and adversarial techniques). 1,800 harmful medical requests violating medical ethics, along with 900 corresponding safe responses. The dataset is structured based on the Principles of Medical Ethics from the American Medical Association (AMA). Assess the medical safety of LLMs by testing whether they refuse to comply with harmful medical requests. Fine-tune LLMs using medical safety demonstrations to enhance their alignment with ethical medical guidelines. Safety
UNIWIZ Synthetic and curated data, including: 17,638 quality-controlled conversations, and 10,000 augmented preference data Features conversations that integrate safety and knowledge alignment. A "safety-priming" method was employed to generate synthetic safety data, and factual information was injected into conversations by retrieving content from curated sources. Fine-tune large language models to enhance their performance in generating safe and knowledge-grounded conversations. Safety
SciFact Curated Domain-Specific Dataset. Includes claims and corresponding evidence abstracts, each annotated with labels indicating whether the claim is supported or refuted, along with rationales justifying the decision. To verify the veracity of scientific claims by identifying supporting or refuting evidence within abstracts and providing justifications for these decisions. Truthfulness
PubHealthTab Curated Domain-Specific Dataset Contains 1,942 real-world public health claims, each paired with evidence tables extracted from over 300 websites. Facilitates evidence-based fact-checking by providing claims and corresponding evidence tables for verification. Truthfulness
LAMA Curated Domain-Specific Dataset. Comprises a set of knowledge sources, each containing a collection of facts. To probe pretrained language models to determine the extent of their factual and commonsense knowledge. Truthfulness
TriviaQA Curated Domain-Specific Dataset. Consists of over 650,000 question-answer pairs, each linked to a set of supporting documents. The questions are sourced from trivia websites, and the answers are derived from the corresponding documents. Training and evaluating models on reading comprehension, specifically focusing on the ability to extract and reason over information from provided documents to answer questions. Truthfulness
Natural Questions (NQ) Real data consists of real anonymized queries from Google’s search engine users, paired with answers derived from entire Wikipedia articles. To develop and evaluate question-answering systems that can read and comprehend entire Wikipedia articles to find answers to user queries. Truthfulness
PopQA Curated Domain-Specific Dataset. consists of 14,000 QA pairs, each associated with fine-grained Wikidata entity IDs, Wikipedia page views, and relationship type information. Designed for open-domain question answering tasks, focusing on evaluating the effectiveness of language models in retrieving and utilizing factual knowledge. Truthfulness
FEVER Curated Domain-Specific Dataset. comprises 185,000 claims, each paired with evidence from Wikipedia articles. These claims are categorized as supported, refuted, or not verifiable. Fact extraction and verification, where models are trained to determine the veracity of claims based on provided evidence. Truthfulness
HEALTHVER Curated Domain-Specific Dataset. Contains 14,330 evidence-claim pairs (SUPPORTS, REFUTES, NEUTRAL) on health claims, mainly COVID-19, verified against scientific articles. Used to train and evaluate models in verifying health claims by classifying them based on scientific evidence. Truthfulness
Datasets Data Type Content Task Dimensions
Med-HALT Synthetic and Real Data, Curated Domain-Specific Dataset, and Public Dataset Consist of Reasoning-Based Assessments, Memory-Based Assessments, Medical Scenarios, Evaluation Metrics Tasks including Evaluation of Hallucination in Medical AI, Reliability Benchmarking, Error Analysis, Mitigation Development Truthfulness
MedICaT Public Text Corpora And Real Data (curated from publicly available biomedical literature) Contains medical images (e.g., radiographs, charts, and diagrams) paired with captions extracted from biomedical literature. Also, includes metadata about the source and context of the images. Task including Medical Image Captioning, Text-Image Retrieval, Medical Reasoning Truthfulness
Med-HallMark Curated Domain-Specific Dataset, , Synthetic and Real Data (includes a mix of real-world medical data and synthetically generated hallucination scenarios) Diverse medical multimodal data, including text, images, and paired annotations. Hierarchically categorized hallucination data, addressing both structural (e.g., object-level) and contextual (e.g., domain knowledge) hallucinations. Task including Hallucination Detection, Hallucination Evaluation, Mitigation Analysis Truthfulness
BioASQ Curated Domain-Specific Dataset; Real Data. The dataset comprises English-language biomedical questions, each accompanied by reference answers and related materials. These questions are designed to reflect real information needs of biomedical experts, making the dataset both realistic and challenging. The primary task is Biomedical Question Answering (QA), which involves systems providing accurate answers to questions based on biomedical data. The dataset supports various QA tasks, including yes/no, factoid, list, and summary questions. Truthfulness
FactualBio Synthetic Data; Public Text Corpora. collection of biographies of individuals notable enough to have Wikipedia pages but lacking extensive detailed coverage. The dataset was generated using GPT-4 and includes biographies of 21 individuals randomly sampled from the WikiBio dataset. Evaluating the factual accuracy of language models, particularly in the context of biography generation. It serves as a benchmark for detecting hallucinations and assessing the factual consistency of generated text. Truthfulness
PubMedQA Curated Domain-Specific Dataset. Consists of over 1,000 question-answer pairs derived from PubMed abstracts, focusing on various biomedical topics. Evaluates the ability of models to comprehend and extract information from biomedical texts to answer specific questions. Truthfulness
MedQuAD Curated Domain-Specific Dataset. The dataset encompasses 37 question types, such as Treatment, Diagnosis, and Side Effects, associated with diseases, drugs, and other medical entities like tests. Designed for medical question answering, the dataset aids in developing and evaluating systems that can understand and respond to medical inquiries. Truthfulness
LiveMedQA2017 Curated Domain-Specific Dataset Consists of 634 question-answer pairs corresponding to National Library of Medicine (NLM) questions Medical question answering, focusing on consumer health questions received by the U.S. National Library of Medicine. Truthfulness
MASH-QA Curated Domain-Specific Dataset. Approximately 25,000 question-answer pairs sourced from WebMD, covering a wide range of healthcare topics. Designed for multiple-answer span extraction in healthcare question answering. Truthfulness
SecureSQL Curated domain-specific dataset Comprises meticulously annotated samples, including both positive and negative instances. The dataset encompasses 57 databases across 34 diverse domains, each associated with specific security conditions. Evaluate and analyze data leakage risks in LLMs, particularly concerning SQL query generation and execution. Privacy
Datasets Data Type Content Task Dimensions
Medical Meadow curated domain-specific dataset It comprises approximately 1.5 million data points across various tasks, including question-answer pairs generated from openly available medical data using models like OpenAI’s Designed to enhance large language models (LLMs) for medical applications Privacy
Electronic Health Records (EHR) at (KHCC) Private dataset gpt-3.5-turbo Clinical research, outcome analysis. Privacy
MedVQA Curated domain-specific dataset A collection of medical visual question answering pairs, designed to train and evaluate models that interpret medical images and answer related questions. Visual question answering, medical image understanding. Explainability
MedExQA Curated domain-specific dataset A dataset focused on medical examination questions and answers, intended to aid in the development of AI models for medical exam preparation and assessment. Question answering, educational assessment. Explainability
MedMCQA Curated domain-specific dataset A multiple-choice question-answering dataset in the medical domain, aimed at training models to handle medical examinations and practice questions. Multiple-choice question answering, medical education. Explainability
TCM Medical Licensing Examination(MLE) Curated domain-specific dataset A dataset comprising questions and answers from Traditional Chinese Medicine licensing examinations. Educational assessment, question answering. Explainability
Pneumonia Dataset Curated domain-specific dataset Medical images (such as chest X-rays) labeled for the presence or absence of pneumonia, used for training diagnostic models. Image classification, disease detection. Explainability
Montgomery Dataset Curated domain-specific dataset Chest X-ray images with manual segmentations of the lung fields, useful for pulmonary research. Image segmentation, tuberculosis detection. Explainability
Shenzhen Dataset Curated domain-specific dataset Chest X-ray images collected in Shenzhen, China, with annotations for tuberculosis manifestations. Disease classification, image analysis. Explainability
IDRID Dataset Curated domain-specific dataset Retinal images with annotations for diabetic retinopathy lesions, intended for retinal image analysis. Image segmentation, disease grading. Explainability
BrainTumor Dataset Curated domain-specific dataset MRI images of brain tumors with corresponding labels, used for developing diagnostic and segmentation models. Tumor detection, image segmentation. Explainability
Table 1: This table provides a structured comparison of datasets used in studies on trust in LLMs for healthcare. The datasets are categorized by data type (e.g., web-scraped, curated domain-specific, synthetic, real-world, or private datasets), content (e.g., medical literature, patient records, clinical guidelines, QA pairs), task (e.g., clinical decision support, medical question-answering, document summarization, biomedical fact-checking, chatbot training), and dimensions of trustworthiness (e.g., truthfulness, privacy, safety, robustness, fairness, bias, explainability). This comparison highlights how each dataset contributes to the development of trustworthy LLMs in medical AI.
Appendix B Comparison of Models
We systematically gathered and analyzed 78 models relevant to studies on trust in LLMs for healthcare. Table 2 provides a comprehensive summary of the LLMs evaluated in these studies, detailing key aspects such as model name, release year, and the institution responsible for its development. Additionally, it specifies the primary task each model is designed for, including medical question-answering, clinical decision support, and biomedical text summarization. To further assess their reliability, we categorize the models based on the dimensions of trustworthiness they address, such as truthfulness, privacy, safety, robustness, fairness and bias, and explainability. This structured overview offers valuable insights into how different LLMs are designed and evaluated to enhance trust in healthcare AI applications.
Models Release Year Institution Primary Task Dimensions
SciBERT 2019 Allen Institute for AI Pre-trained language model specialized for scientific text, particularly biomedical and computer science literature. Fairness and Bias
PaLM-2 2023 Google Multilingual language understanding and generation, with a focus on reasoning and coding tasks. Fairness and Bias
Mixtral-8x70B 2023 Mistral AI Ensemble of language models aimed at improving performance across diverse language tasks. Fairness and Bias, Safety
Med-PaLM 2023 Google Health Specializing in healthcare-related question answering, clinical diagnosis support, and medical literature interpretation. Fairness and Bias
Med-PaLM 2 2024 Google Health Updated version of Med-PaLM, further improving healthcare-related tasks with enhanced accuracy and reliability in medical information retrieval, clinical reasoning, and decision support. Fairness and Bias
Llama-13B 2023 Meta Designed for natural language understanding and generation tasks, such as text summarization, machine translation, and conversational AI. Fairness and Bias
XLNet 2019 Google Research It is used for text classification, question answering, and language modeling tasks. Fairness and Bias
DeBERTa 2020 Microsoft Research Improves BERT and RoBERTa by enhancing the attention mechanism. It performs well in a variety of NLP tasks, such as sentence classification, question answering, and named entity recognition. Fairness and Bias
Llama-7B 2023 Meta Focused on general-purpose natural language understanding and generation, with potential fine-tuning for specific domains like medicine, law, and technology. Fairness and Bias, Truthfulness
Llama 2 70Bchat 2023 Meta Platforms Open-source conversational AI model designed for dialogue and instruction-following tasks. Fairness and Bias, Truthfulness, Safety, Robustness,
GPT-3.5 2022 OpenAI Enhanced language processing capabilities, building upon GPT-3. Fairness and Bias, Truthfulness, Safety, Robustness, Privacy
GPT2 2019 OpenAI Text generation Fairness and Bias, Robustness
PMC Llama 13B 2023 Allen Institute for AI Specialized in medical literature understanding and generation. Fairness and Bias, Robustness
Models Release Year Institution Primary Task Dimensions
GPT-4 2023 OpenAI Advanced language generation and understanding across various domains. Fairness and Bias, Safety, Robustness, Explainability, Privacy
BERT 2018 Google AI Language Pre-trained Transformer model for a wide range of NLP tasks, such as text classification, NER, QA, etc. Fairness and Bias, Safety, Robustness, Truthfulness
LLAMA 2 CHAT 2023 Meta AI Language modeling Robustness, Explainability
MEDALPACA (7B) 2023 medalpaca Medical domain language model fine-tuned for question-answering and medical dialogue tasks. Robustness, Privacy
CLINICAL CAMEL (13B) 2023 the AI and healthcare community Fine-tuned for clinical applications. It is designed to assist with tasks like medical text classification, clinical decision support, information extraction from medical records, and answering clinical questions. Robustness
GPT-2 XL 2019 OpenAI Large-scale language model for text generation and understanding. Robustness
T5-Large 2020 Google Research It treats all NLP tasks as text-to-text tasks, meaning both the input and output are in the form of text, and it’s used for tasks like translation, summarization, and question answering. Robustness
claude-3.5-sonnet 2024 Anthropic It is a variant of Claude, specialized in tasks such as conversational AI, creative writing, poetry generation, and other text-based applications. Robustness
OpenBioLLM-70B 2024 OpenBioAI It is designed to handle tasks such as biological information extraction, gene sequence analysis, protein folding predictions, and other bioinformatics applications. Robustness
BioMistral-7B 2023 Mistral AI Focused on biomedical and healthcare-related text. Its tasks include medical question answering, clinical document analysis, and medical text summarization. Robustness
Medllama3-v20 2024 MedAI Labs Designed to assist in healthcare tasks like clinical reasoning, medical question answering, and patient record analysis. Robustness
ASCLEPIUS (7B) 2023 Asclepius AI Developed for clinical and medical applications, specializing in tasks like diagnosing medical conditions from symptoms, medical text summarization, and extracting structured information from clinical documents. Robustness, Explainability
ALPACA (7B) 2023 Stanford University Fine-tuned version of the LLaMA model aimed at providing high-quality responses to questions, with an emphasis on maintaining ethical and accurate conversational capabilities in diverse domains. Robustness
Models Release Year Institution Primary Task Dimensions
Google’s Bard 2023 Google Conversational AI tool, focused on providing detailed, accurate, and creative responses to user queries. It can handle a variety of tasks, including web search, content generation, and complex QA. Robustness
Text- Davinci-003 2022 OpenAI It is an advanced variant of GPT-3. It is designed for a wide range of natural language understanding and generation tasks, such as answering questions, summarizing text, creative writing, translation, and code generation. Robustness, Truthfulness
LLaMa 2-7B 2023 Meta (formerly Facebook AI Research) Designed to be a general-purpose AI for a wide range of tasks such as text generation, question answering, and summarization, with specific fine-tuning for medical and technical domains. Robustness, Truthfulness, Privacy
ChatGPT 2022 OpenAI Conversational AI Robustness, Truthfulness, Explainability, Privacy
Llama-3.1 2024 Meta AI Multilingual large language model designed for a variety of natural language processing tasks. Safety
ClinicalCamel-70b 2023 the AI and healthcare community Medical language model designed for clinical research applications. Safety, Explainability
Med42-70b 2023 M42 Health Clinical large language model providing high-quality answers to medical questions. Safety, Explainability
GPT-4o 2024 OpenAI Multimodal large language model capable of processing and generating text, audio, and images in real time. Safety, Privacy, Explainability
Mistral 2023 Mistral AI Language model optimized for code generation and reasoning tasks. Safety, Robustness, Explainability
Meditron (7) (70b) 2023 École Polytechnique Fédérale de Lausanne (EPFL) Medical language model fine-tuned for clinical decision support and medical reasoning. Safety, Robustness, Explainability
Claude-2.1 2023 Anthropic General-purpose language model for a wide range of natural language understanding and generation tasks. Safety, Robustness
GPT-J 2021 EleutherAI Open-source language model for text generation and understanding. Safety, Robustness
Vicuna 2023 UC Berkeley and Microsoft Research Conversational AI Safety, Robustness, Truthfulness
Models Release Year Institution Primary Task Dimensions
Medalpaca-13b 2023 medalpaca Medical domain language model fine-tuned for question-answering and medical dialogue tasks. Safety, Truthfulness, Privacy
GPT-3 2020 OpenAI Natural language understanding and generation Truthfulness, Explainability
ALBERT 2019 Google Research Lighter version of BERT that reduces parameters for efficiency while maintaining performance. It excels in tasks such as text classification, named entity recognition, and question answering. Truthfulness
RoBERTa 2019 Facebook AI Research Optimized variant of BERT that removes the Next Sentence Prediction task and trains with more data and for longer periods. It is used for tasks like question answering, sentiment analysis, and text classification. Truthfulness
BlueBERT 2019 NIH and Stanford University BERT-based model pre-trained on clinical and biomedical text. It is designed for healthcare-related tasks, including clinical text classification, named entity recognition, and medical question answering. Truthfulness
ClinicalBERT 2019 University of Pennsylvania Variant of BERT fine-tuned on clinical texts, tailored for clinical NLP tasks like named entity recognition, clinical event extraction, and question answering in the medical domain. Truthfulness
TAPAS 2020 Google Research Designed for answering questions based on tabular data. It is used for tasks like extracting structured information from tables and processing queries in tabular datasets. Truthfulness
LLaMA-2 13B 2023 Meta Advanced variant of Meta’s LLaMA series, designed for text generation, question answering, summarization, and other NLP tasks. Truthfulness, Explainability, Privacy
MPT 2023 MosaicML General-purpose LLM for text generation, summarization, language understanding, and reasoning tasks. Fine-tuned for downstream applications such as chatbot development, code generation, and other NLP tasks. Truthfulness
BLIP2 2023 Salesforce Bootstrapping language-image pre-training, designed to bridge vision-language models with large language models for improved visual understanding and generation. Truthfulness
InstructBLIP-7b/13b 2023 Salesforce Visual instruction-tuned versions of BLIP-2, utilizing Vicuna-7B and Vicuna-13B language models, respectively, to enhance vision-language understanding through instruction tuning. Truthfulness
LLaVA1.5-7b/13b 2023 Microsoft Large language and vision assistant models with 7B and 13B parameters, respectively, designed for multimodal tasks by integrating visual information into language models. Truthfulness
mPLUGOwl2 2023 Zhejiang University Multimodal pre-trained language model designed to handle various vision-language tasks, including image captioning and visual question answering. Truthfulness
Models Release Year Institution Primary Task Dimensions
XrayGPT 2023 University of Toronto Specialized model for generating radiology reports from chest X-ray images, aiming to assist in medical image interpretation. Truthfulness
MiniGPT4 2023 King Abdullah University of Science and Technology A lightweight multimodal model designed to align vision and language models efficiently, facilitating tasks like image captioning and visual question answering. Truthfulness
RadFM 2023 Stanford University Foundation model tailored for radiology, focusing on interpreting medical images and integrating findings with clinical language models. Truthfulness
Alpaca-LoRA 2023 Stanford University It focuses on achieving good performance in tasks such as question answering and personalized dialogue. Truthfulness
Robin- medical 2023 Robin Health Fine-tuned for medical applications, including clinical decision support, medical question answering, and health record analysis. Truthfulness
Flan-T5 2021 Google Research Optimized for tasks like question answering, text summarization, and sentence classification, across a variety of domains. Truthfulness
BioBERT 2019 Korea University Biomedical language representation learning, enhancing performance on tasks like named entity recognition, relation extraction, and question answering within the biomedical domain. Truthfulness
Falcon Instruct (7B and 40B) 2023 Technology Innovation Institute (TII), UAE. Instruction-tuned language model designed to follow user instructions effectively. Truthfulness, Robustness
Mistral Instruct (7B) 2023 Mistral AI Instruction-tuned language model designed to follow user instructions effectively. Truthfulness, Robustness
Falcon 2023 Technology Innovation Institute (TII), UAE. General-purpose language model optimized for text understanding, generation, question answering, and reasoning tasks. Focused on efficient deployment for industry-scale applications. Truthfulness, Robustness
LLaVA-Med 2024 Microsoft Large language and vision assistant for biomedicine, trained to handle visual instruction tasks in the biomedical field, aiming for capabilities similar to GPT-4. Truthfulness, Explainability
GPT-4o-mini 2024 OpenAI Natural language processing (NLP), text generation, and understanding. Explainability
ASCLEPIUS (13B) 2023 Asclepius AI Medical NLP, clinical text analysis, and healthcare-related tasks. Explainability
Models Release Year Institution Primary Task Dimensions
MedViLaM 2023 Xu et al. (2024) Medical vision-language tasks, combining image and text analysis for healthcare. Explainability
Med-MoE 2023 Jiang et al. (2024) Medical NLP, leveraging Mixture of Experts (MoE) for specialized healthcare tasks. Explainability
Gemini Pro 2023 Google DeepMind Multi-modal NLP, combining text, image, and other data types for advanced AI tasks Explainability
AlpaCare (7B) (13B) 2023 Zhang et al. (2023) Healthcare-focused NLP, clinical text analysis, and medical decision support Explainability
Yi (6B) 2023 01.AI (China) General-purpose NLP, text generation, and fine-tuning for specific applications. Explainability
Phi-2 (2.7B) 2023 Microsoft Lightweight NLP, text generation, and fine-tuning for specific tasks. Explainability
SOLAR (10.7B) 2023 Upstage AI General-purpose NLP, text generation, and fine-tuning for specific domains. Explainability
InternLM2 (7B) 2023 Shanghai AI Laboratory (China) General-purpose NLP, text generation, and fine-tuning for specific applications. Explainability
Llama3-( 8B and 70B) 2024 Meta General-purpose NLP, text generation, and fine-tuning for specific applications. Privacy
CodeLlama-( 7B, 13B, and 34B) 2023 Meta Code generation, code completion, and programming assistance. Privacy
Mixtral-8x7B and 8x22B 2023 Mistral AI General-purpose NLP, text generation, and fine-tuning for specific domains. Privacy
Qwen-(7B, 14B, 32B, 72B)-Chat 2023 Alibaba Chat-oriented NLP, conversational AI, and text generation. Privacy
GLM-4 2024 Tsinghua University Advanced NLP, text generation, and multi-modal tasks. Privacy
Table 2: Detailed Comparison of GPT Models Evaluated for Trust in Healthcare LLMs, Including Model Name, Release Year, Institution, Primary Tasks (e.g., Medical Question-Answering, Clinical Decision Support, Biomedical Text Summarization, Medical Report Generation), and Key Trustworthiness Dimensions (Truthfulness, Privacy, Safety, Robustness, Fairness and Bias, Explainability).

### v2

@@
 Abstract
-The application of large language models (LLMs) in healthcare has the potential to revolutionize clinical decision-making, medical research, and patient care.
-As LLMs are increasingly integrated into healthcare systems, several critical challenges must be addressed to ensure their reliable and ethical deployment.
-These challenges include truthfulness, where models generate misleading information; privacy, with risks of unintentional data retention; robustness, requiring defenses against adversarial attacks; fairness, addressing biases in clinical outcomes; explainability, ensuring transparent decision-making; and safety, mitigating risks of misinformation and medical errors.
-Recently, researchers have begun developing benchmarks and evaluation frameworks to systematically assess the trustworthiness of LLMs.
-However, the trustworthiness of LLMs in healthcare remains underexplored, lacking a systematic review that provides a comprehensive understanding and future insights into this area.
-This survey bridges this gap by providing a comprehensive overview of the recent research of existing methodologies and solutions aimed at mitigating the above risks in healthcare.
-By focusing on key trustworthiness dimensions including truthfulness, privacy and safety, robustness, fairness and bias, and explainability, we present a thorough analysis of how these issues impact the reliability and ethical use of LLMs in healthcare.
-This paper highlights ongoing efforts and offers insights into future research directions to ensure the safe and trustworthy deployment of LLMs in healthcare.
+The application of large language models (LLMs) in healthcare holds significant promise for enhancing clinical decision-making, medical research, and patient care.
+However, their integration into real-world clinical settings raises critical concerns around trustworthiness, particularly around dimensions of truthfulness, privacy, safety, robustness, fairness, and explainability.
+These dimensions are essential for ensuring that LLMs generate reliable, unbiased, and ethically sound outputs.
+While researchers have recently begun developing benchmarks and evaluation frameworks to assess LLM trustworthiness, the trustworthiness of LLMs in healthcare remains underexplored, lacking a systematic review that provides a comprehensive understanding and future insights.
+This survey addresses that gap by providing a comprehensive review of current methodologies and solutions aimed at mitigating risks across key trust dimensions.
+We analyze how each dimension affects the reliability and ethical deployment of healthcare LLMs, synthesize ongoing research efforts, and identify critical gaps in existing approaches.
+We also identify emerging challenges posed by evolving paradigms, such as multi-agent collaboration, multi-modal reasoning, and the development of small open-source medical models.
+Our goal is to guide future research toward more trustworthy, transparent, and clinically viable LLMs.
 A Comprehensive Survey on the Trustworthiness of Large Language Models in Healthcare
-Manar Aljohani1, Jun Hou1, Sindhura Kommu1, Xuan Wang1, 1Department of Computer Science, Virginia Tech, Blacksburg, VA, USA (manara, junh, sindhura, xuanw)@vt.edu
+Manar Aljohani $\spadesuit$, Jun Hou $\spadesuit$, Sindhura Kommu $\spadesuit$, Xuan Wang $\spadesuit$ $\spadesuit$ Department of Computer Science, Virginia Tech, Blacksburg, VA, USA {manara, junh, sindhura, xuanw}@vt.edu
+(a) Temporal Trends
+(b) Distribution of Datasets
+(c) Distribution of Models
+Figure 1: Overview of research trends, dataset usage, and model types across key trustworthiness dimensions in healthcare LLM studies: (a) Temporal Trends in Trustworthiness Dimensions Addressed in Medical LLM Studies (2022–2025); (b) Distribution of Dataset Types Across Trustworthiness Dimensions in Healthcare LLM Studies; (c) Distribution of Model Types Across Trustworthiness Dimensions in Healthcare LLM Studies.
 1 Introduction
@@
 It also encompasses the model’s resilience against adversarial attacks, ensuring that external manipulations do not compromise its integrity.
-A truly robust LLM in healthcare must demonstrate stability, reliability, and fairness, even when faced with noisy, ambiguous, or adversarial inputs, thereby safeguarding patient safety and supporting clinical decision-making.
-Similarly, fairness and bias must be addressed to prevent discriminatory patterns in model predictions, which could lead to unequal treatment recommendations and exacerbate healthcare disparities.
+A truly robust LLM in healthcare must demonstrate stability, reliability, and fairness, even when faced with noisy, ambiguous, or adversarial inputs.
+Similarly, fairness and bias must be addressed to prevent discriminatory patterns in model outputs, which could lead to unequal treatment recommendations and exacerbate healthcare disparities.
 Furthermore, the explainability of LLMs, which ensures that model outputs are interpretable and transparent, plays a vital role in fostering trust and allowing informed decision-making by healthcare professionals.
-The lack of transparency in model reasoning complicates clinical adoption and raises concerns about accountability.
+Lack of transparency in model reasoning complicates clinical adoption and raises accountability concerns.
+Clinical deployments of LLMs expose trust gaps across dimensions.
+Med-PaLM and Med-PaLM 2 show truthfulness and safety issues, with hallucinated that could misguide care Singhal et al.
+(2023).
+Integrating LLMs with EHRs in cloud settings risks HIPAA/GDPR violations, prompting on-prem deployment and stronger de-identification Jonnagaddala and Wong (2025).
+Robustness remains problematic; frameworks like MEDIC and CREOLA assess hallucination severity and clinical safety Kanithi et al.
+(2024a); Asgari et al.
+(2025).
+Fairness issues persist, with studies showing that LLMs can perpetuate racial biases in medical recommendations Pfohl et al.
+(2024).
+Finally, explainability challenges were evident in AMIE—a conversational diagnostic agent evaluated in OSCE-style clinical exams—demonstrated strong diagnostic reasoning but lacked transparency compared to human doctors Tu et al.
+(2025).
 Tackling these challenges is essential for the trustworthy and ethical implementation of LLMs in healthcare.
@@
 However, there is currently no systematic review that provides a comprehensive understanding and future insights into this area.
-To bridge this gap, we present a comprehensive survey that explores these trust-related dimensions in detail, reviewing existing datasets, solutions, and methodologies aimed at improving the trustworthiness of LLMs in healthcare.
+To bridge this gap, we present a comprehensive survey that explores these trust-related dimensions in detail, reviewing existing benchmarks and methodologies aimed at improving the trustworthiness of LLMs in healthcare.
 2 Datasets, Models, and Tasks
-We first conducted an extensive search for papers on the trustworthiness of LLMs in healthcare.
-Our search utilized a range of keyword combinations, including terms such as ‘large language models,’ ‘foundation model,’ ‘medical,’ ‘clinical,’ ‘explainability,’ ‘truthfulness,’ ‘trustworthiness,’ ‘safety,’ ‘fairness,’ ‘robustness,’ and ‘privacy.’ We explored several reputable venues, including Arxiv, PubMed, ACL, EMNLP, NAACL, ICML, NeurIPS, ICLR, KDD, Nature, Science, AAAI, and IJCAI, with a focus on recent publications post-2021.
-After reviewing the search results, we identified a total of 30,595 papers.
-Following the removal of duplicates, we narrowed the focus to 69 papers that specifically addressed the truthfulness, privacy, safety, robustness, fairness, bias, and explainability of LLMs in the healthcare domain.
-We then summarized all the datasets, models, and tasks relevant to research on trust in LLMs for healthcare, providing a comprehensive overview of their applications and contributions to this domain.
-The datasets used in studies of trust in LLMs for healthcare are categorized by the dimensions of trustworthiness they address in Appendix A, where we highlight key details such as data type, content, task, and dimensions of trustworthiness.
-The content of each dataset specifies its composition, while the task refers to the primary purpose for which the dataset is utilized.
-The data type varies across datasets and includes web-scraped data, curated domain-specific datasets, public text corpora, synthetic data, real-world data, and private datasets, providing a comprehensive overview of their relevance to healthcare applications.
-The models assessed in studies on trust in LLMs for the healthcare domain are outlined, along with their trustworthiness dimensions, in Appendix B, where we summarized key details such as the model name, release year, task, and the institution responsible for its development.
-The tasks covered various primary focuses of LLMs in healthcare.
-Based on insights from the survey by Liu et al.
-(2024), these tasks are outlined as follows:
+2.1 Inclusion & Exclusion Criteria
+We initiated our survey with a comprehensive literature search targeting studies on the trustworthiness of LLMs in healthcare.
+Our search strategy employed diverse keyword combinations and was directed toward top-tier conferences and journals, prioritizing publications from 2022 onward.
+Detailed inclusion and exclusion criteria are provided in Appendix A.
+Fig 1(a) illustrates how the number of papers addressing each key trustworthiness dimension in healthcare LLMs has changed over time from 2022 to 2025.
+From Figure 1(a), interest in trustworthiness dimensions peaked in 2024, particularly for Fairness and Bias (16 papers) and Privacy (11 papers), reflecting a strong recent push toward ethical and secure AI in healthcare.
+Truthfulness and Explainability maintained steady growth through 2023 and 2024.
+These trends suggest a rising concern with fairness and privacy in recent years, possibly driven by real-world deployment risks and regulatory pressure.
+{forest}
+forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=14em,font=,, where level=2text width=18em,font=,, where level=3text width=18em,font=,, [ Trustworthiness of LLMs in Healthcare, ver [ Truthfulness, fill=blue!10 [ Benchmarks, fill=blue!10 [ Med-HALT Pal et al.
+(2023), PubHealthTab Akhtar et al.
+(2022), HEALTHVER Sarrouti et al.
+(2021), leaf, text width=66em ] ] [ Mitigation Methods, fill=blue!10 [ Self Reflection Ji et al.
+(2023), MEDAL Li et al.
+(2024), Faithful Reasoning Tan et al.
+(2024), HEALTHVER Sarrouti et al.
+(2021),
+CRITIC Gou et al.
+(2024), SEND Mohammadzadeh et al.
+(2024), leaf, text width=66em ] ] [ Evaluation and Detection Methods, fill=blue!10 [ Med-HALT Pal et al.
+(2023), Med-HVL Yan et al.
+(2024), Semantic Entropy Farquhar et al.
+(2024),
+SEPs Han et al. (2024a), Faithful Reasoning Tan et al. (2024), PubHealthTab Akhtar et al. (2022), HEALTHVER Sarrouti et al. (2021),
+CRITIC Gou et al.
+(2024), Cross-Examination Cohen et al.
+(2023), MAD Smit et al.
+(2023), leaf, text width=66em ] ] ] [ Privacy, fill=blue!10 [ Benchmarks, fill=blue!10 [ SecureSQL Song et al.
+(2024), leaf, text width=66em ] ] [ Enhance Methods, fill=blue!10 [ Federated Learning Zhao et al.
+(2024a),
+Differential Privacy Singh et al.
+(2024), De-identification Liu et al.
+(2023b), Mitigating Memorization Yang et al.
+(2024a), APNEAP Wu et al.
+(2024), leaf, text width=66em ] ] [ Evaluation Methods, fill=blue!10 [ SecureSQL Song et al.
+(2024), Memorize Fine-tuning Data Yang et al.
+(2024a), clinical Note De-identification Altalla’ et al.
+(2025),
+Memorization Yang et al.
+(2024a), Textual Data Sanitization Xin et al.
+(2024), leaf, text width=66em ] ] ] [ Safety, fill=blue!10 [ Benchmarks, fill=blue!10 [ Med-harm Han et al.
+(2024c), Medsafetybench Han et al.
+(2024b), leaf, text width=66em ] ] [ Enhance Methods, fill=blue!10 [ UNIWIZ Das and Srihari (2024), Data-Poisoning Attack Han et al.
+(2024e), leaf, text width=66em ] ] [ Evaluation Methods, fill=blue!10 [ Med-harm Han et al.
+(2024c), Medsafetybench Han et al.
+(2024b), Misinformation Attacks Han et al.
+(2024d), MEDIC Kanithi et al.
+(2024b),
+GLiR Attack Leemann et al.
+(2024), Data-Poisoning Attack Han et al.
+(2024e), leaf, text width=66em ] ] ] [ Robustness, fill=blue!10 [ Benchmarks, fill=blue!10 [ Detecting Anomalies Rahman et al.
+(2024), RobustQA Han et al.
+(2023), RABBITS Gallifant et al.
+(2024), leaf, text width=66em ] ] [ Enhance Methods, fill=blue!10 [ LLM-TTA O’Brien et al.
+(2024), Detecting Anomalies Rahman et al.
+(2024), Secure Your Model Tang et al.
+(2024), MEDSAGE Binici et al.
+(2025),
+Out-of-Context Prompting Cotta and Maddison (2024), leaf, text width=66em ] ] [ Evaluation Methods, fill=blue!10 [ Stumbling Blocks Wang et al. (2024), Detecting Anomalies Rahman et al. (2024),
+Instruction Phrasings Ceballos-Arroyo et al.
+(2024), RobustQA Han et al.
+(2023), RABBITS Gallifant et al.
+(2024), leaf, text width=66em ] ] ] [ Fairness and Bias, fill=blue!10 [ Benchmarks, fill=blue!10 [ BiasMedQA Schmidgall et al.
+(2024), EquityMedQA Pfohl et al.
+(2024), Superficial Fairness Alignment Wei et al.
+(2024), FairMedFM Jin et al.
+(2024), leaf, text width=66em ] ] [ Mitigation Methods, fill=blue!10 [ BiasMedQA Schmidgall et al.
+(2024), Reinforcement Learning with Clinician Feedback Zack et al.
+(2024), Instruction Fine-tuning Singhal et al.
+(2023),
+Hurtful Words Zhang et al. (2020), Mitigate Cognitive Biases Ke et al. (2024), CI4MRC Zhu et al. (2023), Bias of Disease Prediction Zhao et al. (2024b),
+Racial and LGBTQ+ Biases Xie et al. (2024), Out-of-Context Prompting Cotta and Maddison (2024), Attribute Neutral Modeling Hu et al. (2024),
+Personalized Alignment Techniques Kirk et al.
+(2024), Evaluating Biases in Context-Dependent Patel et al.
+(2024), leaf, text width=66em ] ] [ Evaluation and Detection Methods, fill=blue!10 [ Evaluation Study Zack et al.
+(2024), BiasMedQA Schmidgall et al.
+(2024), Hurtful Words Zhang et al.
+(2020), Race-based Medicine Omiye et al.
+(2023),
+Detect Debunked Stereotypes Swaminathan et al. (2024), EquityMedQA Pfohl et al. (2024), Superficial Fairness Alignment Wei et al. (2024),
+Examines Biased AI Adam et al. (2022), Identify Biases Yang et al. (2024b), Quantifying Cognitive Biases Lin and Ng (2023),
+Biases in Biomedical MLM Kim et al. (2023), Bias of Disease Prediction Zhao et al. (2024b), Racial and LGBTQ+ Biases Xie et al. (2024),
+FairMedFM Jin et al.
+(2024), leaf, text width=66em ] ] ] [ Explanability, fill=blue!10 [ Benchmarks, fill=blue!10 [ FaReBio Fang et al.
+(2024), Pathway2Text Yang et al.
+(2022), leaf, text width=66em ] ] [ Enhance Methods, fill=blue!10 [ Knowledge Graphs Shariatmadari et al.
+(2024), Medical Imaging Explainability Ghosh et al.
+(2023), MedExQA Kim et al.
+(2024),
+Retrieval and Reasoning on KGs Ji et al. (2024), DDCoT Zheng et al. (2023), A ChatGPT Aided Explainable Framework Liu et al. (2023a),
+Medical Concept-Driven Attention Wang et al. (2022), FaReBio Fang et al. (2024),
+LLM-GCE He et al. (2024), kNN-Graph2Text Yang et al. (2022), RAG-IM Mahbub et al. (2024), MedThink Gai et al. (2025), leaf, text width=66em ] ] ] ]
+Figure 2: Summary of the recent research across various dimensions of trustworthiness of LLMs in healthcare.
+2.2 Datasets
+The datasets used in studies of trust in LLMs for healthcare are categorized by the dimensions of trustworthiness they address in Appendix B, where we highlight key details such as data type, content, task, and dimensions of trustworthiness.
+The content of each dataset specifies its composition, while the task refers to the main problem to be solved for which the dataset is utilized.
+The data type varies across studies and includes web-scraped data, curated domain-specific datasets, public text corpora, synthetic data, real-world data, and private datasets, providing a comprehensive overview of their relevance to healthcare applications.
+Figure 1(b) shows the number of studies using three major dataset types—Med-QA (blue), Med-Gen (orange), and Med-IE (green)—in relation to six trustworthiness dimensions: Truthfulness, Privacy, Safety, Robustness, Fairness and Bias, and Explainability.
+Figure 1(b) shows how three major dataset types—Med-QA, Med-Gen, and Med-IE—are used across six trust dimensions.
+Truthfulness is most studied with both Med-QA and Med-Gen.
+Med-QA is also common in fairness and explainability, while Med-Gen contributes to safety and privacy.
+Med-IE, though less used overall, is more prominent in robustness and explainability.
+This highlights the dominance of Med-QA and Med-Gen, with Med-IE offering value in specific areas of trustworthiness.
+2.3 Models
+The models assessed in studies on trust in LLMs for the healthcare domain are outlined, along with their trustworthiness dimensions, in Appendix C, where we summarized key details such as the model name, release year, openness, architecture, task, and the institution responsible for its development.
+Figure 1(c) illustrates the proportions of different model types—open-source, closed-source, and architectures including encoder-only, decoder-only, and encoder-decoder—used in research addressing various trustworthiness aspects of LLMs in healthcare: Explainability, Fairness and Bias, Robustness, Safety, Privacy, and Truthfulness.
+From Figure 1(c), it is clear that Decoder-only and Open-source models are the most commonly used across all trustworthiness dimensions—especially in robustness, explainability, and truthfulness—highlighting their accessibility and alignment with generative tasks.
+Closed-source models appear more in fairness and privacy studies, while Encoder-only and Encoder-decoder models are used less frequently, mostly in fairness and truthfulness evaluations.
+2.4 Tasks
+The tasks covered various primary focuses of LLMs in healthcare. Inspired from the survey by Liu et al. (2024a), these tasks include:
 Medical Information Extraction (Med-IE)
@@
 Their core functions include query understanding (interpreting user questions), information retrieval (finding relevant data in medical databases), and inference and reasoning (drawing conclusions, inferring relationships, and predicting outcomes based on retrieved data).
-{forest}
-forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=14em,font=,, where level=2text width=18em,font=,, where level=3text width=18em,font=,, [ Trustworthiness of LLMs in Healthcare, ver [ Truthfulness, fill=blue!10 [ Med-HALT Pal et al.
-(2023), Med-HVL Yan et al.
-(2024b), Med-HallMark Chen et al.
-(2024), Semantic Entropy Farquhar et al.
-(2024),
-SEPs Han et al. (2024a), Survey Ahmad et al. (2023), Self Reflection Ji et al. (2023), Model-Agnostic Hallucination Post-Processing Li et al. (2024),
-Faithful Reasoning Tan et al. (2024), PubHealthTab Akhtar et al. (2022), HEALTHVER Sarrouti et al. (2021), CRITIC Gou et al. (2024),
-Cross-Examination Cohen et al.
-(2023), KnowledgeEditor De Cao et al.
-(2021), leaf, text width=66em ] ] [ Privacy, fill=blue!10 [ Survey Das et al.
-(2024), Survey Yan et al.
-(2024a), Privacy Risks Pan et al.
-(2020), Federated Learning Zhao et al.
-(2024),
-Differential Privacy Singh et al. (2024), SecureSQL Song et al. (2024), Memorize Fine-tuning Data Yang et al. (2024a),
-clinical Note De-identification Altalla’ et al.
-(2025), De-identification Liu et al.
-(2023b), leaf, text width=66em ] ] [ Safety, fill=blue!10 [ Med-harm Han et al.
-(2024c), Medsafetybench Han et al.
-(2024b), UNIWIZ Das and Srihari (2024), Misinformation Attacks Han et al.
-(2024d),
-MEDIC Kanithi et al.
-(2024), leaf, text width=66em ] ] [ Robustness, fill=blue!10 [ Survey Yuan et al.
-(2023), Survey Alberts et al.
-(2023), Stumbling Blocks Wang et al.
-(2024), LLM-TTA O’Brien et al.
-(2024), MedFuzz Ness et al.
-(2024),
-Detecting Anomalies Rahman et al. (2024), Instruction Phrasings Ceballos-Arroyo et al. (2024), Adversarial Attacks Yang et al. (2024b),
-Secure Your Model Tang et al.
-(2024), leaf, text width=66em ] ] [ Fairness and Bias, fill=blue!10 [ Evaluation StudyZack et al.
-(2024), BiasMedQA Schmidgall et al.
-(2024), Survey Parray et al.
-(2023), Instruction Fine-tuning Singhal et al.
-(2023),
-Hurtful Words Zhang et al. (2020), Race-based Medicine Omiye et al. (2023), Detect Debunked Stereotypes Swaminathan et al. (2024),
-EquityMedQA Pfohl et al. (2024), Superficial Fairness Alignment Wei et al. (2024), Examines Biased AI Adam et al. (2022),
-Identify Biases Yang et al.
-(2024c), Quantifying Cognitive Biases Lin and Ng (2023), Mitigate Cognitive Biases Ke et al.
-(2024), CI4MRC Zhu et al.
-(2023), leaf, text width=66em ] ] [ Explanability, fill=blue!10 [ Knowledge Graphs Shariatmadari et al.
-(2024), LLMs and Explainable ML Elsborg and Salvatore (2023), Medical Imaging Explainability Ghosh et al.
-(2023),
-MedViLaM Xu et al. (2024), MedThink Gai et al. (2024), MedExQA Kim et al. (2024), Causal Graphs Meet Thoughts Luo et al. (2025),
-Retrieval and Reasoning on KGs Ji et al. (2024), TOSRR Liu et al. (2025), DDCoT Zheng et al. (2023), Layered Chain-of-Thought Prompting Sanwal (2025),
-A ChatGPT Aided Explainable Framework Liu et al. (2023a), leaf, text width=66em ] ] ]
-Figure 1: Summary of the recent research across various dimensions of trustworthiness of LLMs in healthcare.
 Medical Natural Language Inference (Med-NLI)
@@
 We examine the challenges related to the trustworthiness of LLMs in healthcare, outlining key strategies for identifying and mitigating these concerns.
-From our literature review screening, we identified truthfulness, privacy and safety, robustness, fairness and bias, and explainability as key trustworthiness dimensions of LLMs as highlighted in TrustLLM Huang et al.
+From our literature review screening, we identified truthfulness, privacy, safety, robustness, fairness and bias, and explainability as key trustworthiness dimensions of LLMs as highlighted in TrustLLM Huang et al.
 (2024), particularly in healthcare.
@@
 3.1 Truthfulness
-Ensuring the truthfulness of LLMs in healthcare is vital, as inaccurate information can significantly impact patient care and clinical outcomes.
-Given their influence on diagnoses and treatment decisions, it is essential to develop effective methods to detect and mitigate hallucinations and factual inaccuracies.
-Hallucinations in medical LLMs arise from reliance on unverified sources, biases in training data, and limitations in contextual understanding and sequential reasoning Ahmad et al.
+Findings in Truthfulness Self-reflection and fact-checking reduce hallucinations but do not scale or generalize well, especially for long-form clinical contexts.
+Improving truthfulness will require hybrid pipelines that combine retrieval, reasoning, multi-agent self-correction, and uncertainty estimation.
+Ensuring the truthfulness of LLMs in healthcare is critical, as inaccurate or fabricated information can directly harm clinical decisions.
+Hallucinations arise from biased data, weak contextual reasoning, and unverifiable sources Ahmad et al.
 (2023).
-Addressing these issues requires robust evaluation frameworks, self-correction mechanisms, and uncertainty quantification techniques.
+Current work targets factual reliability via benchmarking, post-hoc correction, uncertainty quantification, and improved evidence synthesis.
+Several benchmarks have emerged to quantify and categorize hallucinations.
 The Med-HALT benchmark Pal et al.
-(2023) is designed to evaluate hallucinations in medical LLMs by using reasoning-based tests like ‘False Confidence’ and ‘None of the Above,’ as well as memory-based tests to assess how well the model recalls medical knowledge.
-On the other hand, the interactive self-reflection methodology Ji et al.
-(2023) aims to reduce hallucinations in medical question-answering tasks by introducing an iterative feedback loop where the model refines its responses through self-evaluation and knowledge adjustment.
-In the context of multimodal models, Med-HVL Yan et al.
-(2024b) introduces two key metrics—Object Hallucination and Domain Knowledge Hallucination—to quantify hallucinations in Large Vision-Language Models (LVLMs).
-This framework also uses the CHAIR (Caption Hallucination Assessment with Image Relevance) metric to assess object hallucinations in image captioning.
-Med-HallMark Chen et al.
-(2024) takes it a step further by providing a multi-task evaluation framework with a hierarchical categorization of hallucinations, introducing the MediHall Score for assessing hallucination severity and the MediHallDetector, a multitask-trained LVLM for hallucination detection.
-Researchers have also investigated semantic entropy, a probabilistic measure of uncertainty, to detect hallucinations in LLMs.
-For example, Farquhar et al.
-(2024) leverage semantic entropy to identify confabulations—hallucinations where the model generates arbitrary or incorrect outputs.
-While effective, this approach is computationally expensive, limiting its scalability.
-To overcome this, Han et al.
-(2024a) introduce Semantic Entropy Probes (SEPs), which approximate semantic entropy directly from hidden states.
-By eliminating the need for multiple output samples, SEPs significantly reduce computational overhead.
-Both methods have been successfully applied to biomedical datasets, such as BioASQ.
-Although these techniques offer valuable contributions, hallucination mitigation methods often lack adaptability, being either task-specific or requiring expensive retraining.
-To address this gap, MEDAL Li et al.
-(2024) introduces a model-agnostic post-processing framework that integrates with any medical summarization model.
-MEDAL uses a self-examining correction model to improve factual accuracy without adding extra computational costs, providing a practical solution to the issue of medical hallucinations.
-Collectively, these studies underscore the multifaceted challenge of ensuring truthfulness in medical LLMs.
-By leveraging benchmarking frameworks, self-correction mechanisms, and entropy-based uncertainty measures, researchers can develop complementary strategies for detecting, mitigating, and quantifying hallucinations.
-A key focus of these efforts is the development of quantifiable scoring methods, enabling systematic assessment and comparison across different models.
-These evaluation techniques not only help identify the most reliable and effective LLMs for healthcare applications but also provide actionable insights for further improvements.
-Factual accuracy is fundamental to building trust in LLMs, especially in healthcare, where reliable and verifiable information is critical.
-However, current LLMs lack effective mechanisms to trace claims back to their original sources, underscoring the urgent need for improved validation techniques to ensure safe and trustworthy medical applications.
-To address these challenges, several studies have introduced innovative approaches to enhance the transparency, accuracy, and reliability of healthcare LLMs.
-Tan et al.
-(2024) propose an approach that integrates multiple perspectives from scientific literature to evaluate conflicting arguments, thereby improving LLM reasoning.
-Similarly, Akhtar et al.
-(2022) introduce PubHealthTab, a table-based dataset designed for validating public health claims against noisy evidence, while Sarrouti et al.
-(2021) present HEALTHVER, a dataset tailored for evidence-based fact-checking of health-related claims.
-These structured benchmarks provide a foundation for assessing and refining the reliability of LLM-generated medical information.
-Beyond dataset-driven validation, self-correction mechanisms have been explored to improve LLM truthfulness.
+(2023) evaluates hallucination types using reasoning-based tests (e.g., “False Confidence”) and memory checks.
+In multimodal settings, Med-HVL Yan et al.
+(2024) distinguishes between Object Hallucination and Domain Knowledge Hallucination.
+To mitigate hallucinations, post-hoc correction techniques are gaining traction.
+MEDAL Li et al.
+(2024) presents a model-agnostic self-correction module that improves summarization outputs without retraining.
+Similarly, interactive feedback strategies like self-reflection loops Ji et al.
+(2023) allow LLMs to iteratively refine their responses.
+Uncertainty quantification approaches provide complementary detection tools.
+Farquhar et al.
+(2024) apply semantic entropy to flag low-confidence responses, while SEPs Han et al.
+(2024a) offer a lightweight, hidden-state-based approximation suited for clinical use.
+Recent efforts also examine the trustworthiness of evidence synthesis pipelines.
+Zhang et al.
+(2024) highlight risks when LLMs generate clinical summaries without grounding, emphasizing the need for transparency in literature retrieval and evidence aggregation.
+Debate-based evaluation, as explored in MAD Smit et al.
+(2023), introduces multi-agent deliberation to vet factual consistency in medical QA.
+Finally, SEND Mohammadzadeh et al.
+(2024) introduces a neuron dropout technique to detoxify hallucination-prone neurons during training, aiming to improve inherent model truthfulness.
+Factual accuracy is critical for trust in healthcare LLMs, where clinical safety relies on reliable, verifiable outputs.
+Yet, current models often produce ungrounded content and lack source traceability.
+Recent work addresses this through medical claim benchmarks, self-correction, automated fact-checking, multi-turn verification, and multi-perspective reasoning—advancing transparency, factuality, and clinical relevance.
+To support systematic validation, Akhtar et al.
+(2022) introduce PubHealthTab, a table-based dataset for checking public health claims against noisy evidence, while Sarrouti et al.
+(2021) propose HEALTHVER, a benchmark for evidence-based fact-checking tailored to medical claims.
+These resources enable structured evaluation of LLM outputs and form the foundation for improving medical claim verification.
+Beyond static benchmarks, dynamic self-correction methods have shown promise.
 Gou et al.
-(2024) introduce CRITIC, a framework inspired by human fact-checking practices, enabling LLMs to validate and refine their responses through iterative feedback and evaluation.
-Expanding on automated fact-checking, Cohen et al.
-(2023) propose a cross-examination framework, where an examiner LLM identifies inconsistencies through multi-turn interactions with the original model.
-Unlike fully automated verification pipelines, CRITIC incorporates human-like evaluation strategies, enhancing the trustworthiness of fact-checking in medical contexts.
-Overall, these studies advance the factual accuracy and transparency of LLMs in healthcare by introducing structured benchmarks, iterative validation processes, and automated fact-checking strategies.
-By incorporating these approaches, researchers can enhance the reliability of medical LLMs, ensuring they deliver more accurate, evidence-based insights to support clinical decision-making.
+(2024) propose CRITIC, a framework inspired by human fact-checking, in which LLMs iteratively assess and revise their own responses.
+This process mimics expert reasoning and introduces a layer of critical reflection into model outputs.
+Complementing this, Cohen et al.
+(2023) present a cross-examination approach, where a second "examiner" model engages in multi-turn dialogue to probe for factual inconsistencies in the original response.
+While CRITIC emphasizes human-like evaluation, cross-examination leverages interaction between models to simulate external verification.
+To further reduce hallucinations and improve factual consistency, Tan et al.
+(2024) introduce a method that incorporates multiple scientific perspectives when resolving conflicting arguments, strengthening LLMs’ reasoning capabilities through broader contextual understanding.
+Evaluations
+Truthfulness is assessed with hallucination/factuality benchmarks (e.g., Med-HALT) and feedback-loop strategies; expert annotations on HealthSearchQA, MedQA, and MultiMedQA are common but costly and subjective.
+Limitations
+Evaluations remain fragmented: narrow task coverage, varying definitions, closed-source dependencies, and limited generalizability across modalities/clinical domains.
+Many self-correction methods are task-specific and lack robustness.
 3.2 Privacy
-LLM-based healthcare applications pose significant privacy risks due to their ability to memorize and reproduce sensitive patient data Das et al.
+Findings in Privacy LLMs pose serious privacy risks from memorizing and regenerating PHI.
+Differential privacy and federated learning help but often hurt utility; future work needs fine-grained, instance-level risk estimation across training and inference.
+LLMs in healthcare face end-to-end privacy risks due to their tendency to memorize and potentially regenerate sensitive data such as protected health information (PHI) Das et al.
+(2024); Pan et al.
 (2020).
-Unauthorized data exposure can lead to confidentiality breaches, ethical concerns, and compliance violations Pan et al.
-(2020).
-Addressing these risks requires privacy safeguards at different stages of model development.
-A major challenge is unintended data retention and leakage, where LLMs memorize fine-tuning data, increasing re-identification risks.
-Studies show that domain-specific LLMs, such as Medalpaca, can retain sensitive data, making privacy breaches more likely Yang et al.
+Key threats include data memorization, insufficient de-identification, and the privacy-utility trade-offs of fine-tuning methods.
+This section examines current vulnerabilities, mitigation strategies, and emerging approaches for achieving privacy-preserving healthcare LLMs.
+Data memorization is a core concern, especially in domain-specific models like Medalpaca Han et al.
+(2025), which are more likely to retain PHI and pose heightened re-identification risks Yang et al.
 (2024a).
-Additionally, adversarial attacks like prompt injection and inference attacks can further exploit these vulnerabilities, as demonstrated by the SecureSQL benchmark Song et al.
-(2024).
-To mitigate these risks, pre-training privacy safeguards focus on de-identification.
-Altalla’ et al.
-(2025) assess GPT-3.5 and GPT-4 in clinical note de-identification and synthetic data generation.
-Similarly, Liu et al.
-(2023b) propose a GPT-4-enabled framework for masking private information while maintaining text structure.
-However, de-identification remains imperfect, as attackers may infer sensitive details from anonymized text.
-During fine-tuning, techniques such as federated learning Zhao et al.
-(2024) and differential privacy Singh et al.
-(2024), as highlighted by Liu et al.
-(2024), play a crucial role in safeguarding patient data.
-Federated learning enables decentralized training without sharing raw data, but it demands high computational resources.
-Differential privacy adds noise to protect sensitive information but can reduce model accuracy.
-Adversarial defenses remain limited.
-The SecureSQL benchmark Song et al.
-(2024) highlights LLM vulnerabilities to structured query attacks.
-While chain-of-thought (COT) prompting offers partial mitigation, it does not eliminate the risk of data exposure.
-Researchers address privacy concerns in healthcare LLMs through two primary approaches: de-identification, which alters real data to prevent re-identification, and synthetic data generation, which creates artificial data to eliminate reliance on sensitive patient information.
-While these strategies enhance privacy protection and maintain model effectiveness, challenges remain in long-term memorization control and adversarial robustness, requiring further research to strengthen data security and prevent unintended information retention.
+Structured attacks like those demonstrated in SecureSQL Song et al.
+(2024) reveal that even chain-of-thought (CoT) prompting provides only marginal defense against leakage.
+Pre-training privacy measures include de-identification techniques like GPT-4 masking Liu et al.
+(2023b) and synthetic note generation Altalla’ et al.
+(2025), though these offer limited protection.
+Xin et al.
+(2024) caution that such methods may create a false sense of security, as subtle semantic cues can still lead to PHI leakage.
+Fine-tuning methods such as federated learning Zhao et al.
+(2024a) and differential privacy (DP) Singh et al.
+(2024) provide stronger safeguards by decentralizing data or adding noise to protect individual records.
+However, these methods often compromise model performance or scalability Liu et al.
+(2024a).
+Emerging techniques seek to reduce this trade-off.
+APNEAP Wu et al.
+(2024) introduces activation patching for privacy neuron editing, reducing leakage without harming utility.
+Complementarily, Chen and Esmaeilzadeh (2024) offer a broader survey of privacy risks and solutions across generative AI use cases in healthcare.
+Ethical and personalization challenges further complicate privacy design.
+Zhui et al.
+(2024) emphasize building privacy-conscious frameworks in medical education, while Kirk et al.
+(2024) caution that overly personalized alignment strategies may inadvertently violate user privacy, advocating instead for bounded personalization.
+Evaluations
+Typical assessments use membership/attribute inference and reconstruction attacks, plus privacy–utility analyses (e.g., redaction or DP) under controlled settings.
+Real-world audits are scarce, and proposed risk-prediction or memorization-severity frameworks lack medical-specific benchmarks.
+Limitations
+Current defenses (e.g., DP, redaction) still trade performance for privacy.
+Many evaluations assume idealized adversaries, and systematic ways to balance memorization risk and utility—especially for multimodal, long-context models—remain limited.
 3.3 Safety
-Ensuring the safety of LLMs in healthcare is critical, as these models must not generate harmful responses.
-A key safety concern, noted by Han et al.
-(2024d), is that modifying just 1.1 % of a model’s weights can embed persistent biomedical inaccuracies while maintaining overall performance.
-This highlights the need for rigorous validation mechanisms and safety assessments to prevent misleading medical information before clinical use.
-To systematically assess safety, MedSafetyBench Han et al.
-(2024b) was introduced as the first benchmark designed to evaluate LLM safety in medical contexts.
-It includes 1,800 harmful medical queries alongside safety-optimized responses generated using advanced LLMs and adversarial techniques.
-Results indicate that publicly available medical LLMs fail to meet safety standards, but fine-tuning with MedSafetyBench significantly improves safety without compromising performance.
-A major challenge is that adversarial actors can manipulate LLMs to generate unsafe outputs, while excessive safety alignment may induce hallucinations.
-To address this, Das and Srihari (2024) propose UNIWIZ, a two-step framework that unifies safety alignment and factual knowledge retrieval.
-Their safety-priming approach synthesizes safety-focused training data, while a retrieval mechanism ensures that model outputs remain factually accurate.
-Models fine-tuned on UNIWIZ outperform larger state-of-the-art instruction-tuned models across multiple safety and accuracy metrics.
-Another key contribution to safety alignment is from Han et al.
-(2024c), who provide the first comprehensive safety evaluation for medical LLMs (MedLLMs).
-They define key concepts of medical safety and alignment and introduce Med-Harm, a dataset designed to evaluate both general and medical-specific risks.
-This dataset assesses how well LLMs handle harmful medical questions, ensuring they adhere to safety and ethical standards in medical AI.
-Further advancing safety assessments, Kanithi et al.
-(2024) introduce MEDIC, a multi-dimensional trustworthiness evaluation framework that systematically assesses medical LLMs across critical dimensions of clinical competence including medical reasoning and clinical safety.
-These studies collectively offer a multi-faceted approach to LLM safety in healthcare.
-MedSafetyBench Han et al.
-(2024b) provides a standardized benchmark for safety evaluation and fine-tuning, while UNIWIZ Das and Srihari (2024) introduces a structured framework that prevents hallucinations while reinforcing safety.
-Han et al.
-(2024c) focus on comprehensive safety alignment, establishing Med-Harm to evaluate domain-specific risks.
-Lastly, MEDIC Kanithi et al.
-(2024) offers a holistic evaluation franmework for improving practical application of LLMs in clinical settings.
-Together, these efforts contribute to a more rigorous and systematic framework for assessing and improving LLM safety in medical applications.
+Findings in Safety Medical LLMs can still produce harmful or misleading content after safety tuning.
+Benchmarks expose vulnerabilities to adversarial prompts and embedded misinformation.
+Robust safety demands proactive alignment and multi-stage, clinically grounded evaluations that simulate realistic misuse.
+Ensuring safety is critical: small weight edits ($\sim 1.1\%$) can implant lasting biomedical falsehoods without hurting average performance Han et al.
+(2024d), and poisoning as little as 0.001% of training data can embed persistent misinformation Han et al.
+(2024e).
+Key concerns include the ease of injecting persistent falsehoods into model weights, inadequate performance on harmful prompts, trade-offs between safety alignment and hallucination, and privacy-related vulnerabilities that can escalate safety risks.
+This section explores current benchmarks, safety alignment strategies, and the overlap between safety and privacy threats.
+To systematically evaluate harmful outputs, benchmarks like MedSafetyBench Han et al.
+(2024b) and Med-Harm Han et al.
+(2024c) use adversarial and real-world queries to test model responses.
+Results show that even medically fine-tuned LLMs often fail safety criteria unless specifically optimized.
+MEDIC Kanithi et al.
+(2024b) broadens this evaluation across dimensions such as reasoning and reliability, offering a holistic safety diagnostic tool.
+Safety alignment remains challenging due to its tension with other objectives.
+UNIWIZ couples safety-driven training with retrieval to reduce unsafe outputs while preserving accuracy Das and Srihari (2024).
+However, over-alignment increases hallucination, whereas under-alignment permits unsafe behavior, demonstrating the delicate balance required for clinical reliability.
+Finally, privacy threats intersect with safety risks.
+Leemann et al.
+(2024) show that membership inference attacks, like Gradient Likelihood Ratio (GLiR), can detect whether individual patient data was used in training.
+This not only violates privacy but also raises safety concerns, as misuse of sensitive information can misguide clinical outcomes.
+Evaluations
+Methods include manual red teaming, automated stress tests, and healthcare-specific tasks (e.g., drug interactions, diagnostic advice) with expert review, though many prompts still derive from general domains.
+Limitations
+Mitigations often target generic harms rather than medical-specific risks (e.g., dangerous dosing).
+Red teaming rarely involves medical experts, and clinically grounded benchmarks with expert-in-the-loop validation remain limited.
 3.4 Robustness
-Enhancing the robustness of LLMs is crucial for their reliability in healthcare applications.
-A key approach involves developing adversarial test samples tailored to the medical domain, such as synthetic anomaly cases Yuan et al.
-(2023) and boundary stress testing Wang et al.
-(2024), to assess model resilience.
-However, creating clinically meaningful adversarial samples presents unique challenges, as Alberts et al.
-(2023) highlight the need to align adversarial testing methods with the complexities of real-world medical data, where medical dependencies must be accounted for.
-In addition to adversarial testing, uncertainty quantification is another important avenue to improve robustness.
+Findings in Robustness LLMs are fragile under distribution shifts, adversarial prompts, and instruction changes.
+Existing defenses (adversarial testing, test-time adaptation) are often task-specific.
+Robustness demands context-aware evaluation, multi-agent training, and resilience to real-world perturbations.
+Ensuring robustness is vital for clinical deployment.
+Challenges include adversarial vulnerability, sensitivity to domain shifts and instruction variations, and prompt-based attacks.
+To address these issues, recent work explores adversarial testing, test-time adaptation, prompt security, data augmentation, and instruction robustness strategies.
+Adversarial robustness is addressed through synthetic data generation.
+Yuan et al.
+(2023) and Wang et al.
+(2024) introduce adversarial test samples tailored to the medical domain, such as synthetic anomaly cases and boundary stress testing, to assess model resilience.
+Alberts et al.
+(2023) emphasize the importance of aligning adversarial testing methods with real-world medical complexities.
+In parallel, Gallifant et al.
+(2024) reveal that simply substituting generic and brand drug names within biomedical benchmarks leads to performance drops of up to 10%, highlighting the fragility of LLMs to clinically trivial lexical shifts.
+Uncertainty quantification offers another avenue for robustness.
 LLM-TTA O’Brien et al.
-(2024) explores test-time adaptation techniques to enhance model performance on rare or unfamiliar cases, which are common in medical diagnostics.
-Unlike adversarial robustness, which focuses on resistance to manipulated inputs, uncertainty quantification aims to identify when models are likely to be incorrect, providing a complementary safety mechanism.
-Another critical question is whether benchmark performance truly reflects medical robustness.
-MedFuzz Ness et al.
-(2024) challenges assumptions in MedQA by modifying questions to test if models rely on rigid, dataset-specific patterns rather than genuine clinical reasoning.
-This research exposes vulnerabilities in LLMs, revealing that subtle changes in input can significantly impact performance, raising concerns about their reliability in dynamic medical environments.
-Instruction robustness is also a growing concern.
-Ceballos-Arroyo et al.
-(2024) examine how variations in medical instructions affect performance across different LLMs, finding that specialized medical models may be more fragile than general-purpose models when instructions are reworded.
-This counterintuitive result suggests that excessive domain adaptation may decrease flexibility and reduce robustness.
-Adversarial vulnerabilities also pose direct security risks.
-Yang et al.
-(2024b) investigate two adversarial attack strategies across medical tasks using real patient data, demonstrating that fine-tuned models are especially vulnerable to poisoning attacks that subtly alter learned weights.
-While adversarial data does not always degrade general performance, it can introduce dangerous biases into specific medical predictions, making early attack detection a priority.
-To protect against adversarial manipulations, Tang et al.
-(2024) introduce Secure Your Model, a framework that strengthens LLM robustness with cryptographic prompt authentication.
-This mechanism ensures that only verified and secure prompts are processed, mitigating vulnerabilities associated with prompt injections and adversarial attacks, and reducing the risk of model exploitation in medical contexts.
-While all these studies address LLM robustness, they differ in their primary focus.
-MedFuzz Ness et al.
-(2024) and Ceballos-Arroyo et al.
-(2024) expose vulnerabilities in existing benchmarks and task instructions, questioning whether current evaluation methods truly measure robustness or merely reflect dataset biases.
-In contrast, Yang et al.
-(2024b) and Alberts et al.
-(2023) highlight adversarial threats, demonstrating how medical LLMs can be subtly manipulated, raising concerns about their security in real-world applications.
-Meanwhile, LLM-TTA O’Brien et al.
-(2024) takes a different approach, focusing on uncertainty quantification rather than adversarial resistance to enhance reliability in handling unfamiliar cases.
-Secure Your Model Tang et al.
-(2024) provides an additional security layer by introducing proactive adversarial defenses through prompt protection mechanisms, ensuring resilience against manipulation risks.
-These studies highlight that robustness is a multifaceted challenge, requiring advancements in evaluation methods and defensive mechanisms.
-Ensuring LLMs can handle adversarial scenarios, integrate domain knowledge, and adapt to language variations is crucial for their safe deployment in healthcare.
-Strengthening robustness through testing and resilience-building enhances the trustworthiness of medical LLMs, making them more reliable in complex clinical settings.
+(2024) explores test-time adaptation techniques to enhance model performance on rare or unfamiliar cases, common in medical diagnostics.
+This approach complements adversarial robustness by identifying instances where models are likely to err.
+Instruction robustness is examined by Ceballos-Arroyo et al.
+(2024), who find that specialized medical models may be more fragile than general-purpose models when instructions are reworded, suggesting that excessive domain adaptation may reduce flexibility.
+Prompt security is enhanced by Tang et al.
+(2024), who introduce a framework that strengthens LLM robustness with cryptographic prompt authentication, mitigating vulnerabilities associated with prompt injections and adversarial attacks.
+Data augmentation techniques are employed in MEDSAGE Binici et al.
+(2025), which uses LLM-generated synthetic dialogues to simulate ASR errors, improving the robustness of medical dialogue summarization systems.
+Similarly, RobustQA Han et al.
+(2023) benchmarks the robustness of domain adaptation for open-domain question answering across diverse domains, facilitating the evaluation of ODQA’s domain robustness.
+Lastly, prompt engineering strategies, such as out-of-context prompting, are explored by Cotta and Maddison (2024), who demonstrate that applying random counterfactual transformations can improve the fairness and robustness of LLM predictions without additional data or fine-tuning.
+Evaluations
+Robustness is tested via distribution shifts, adversarial or out-of-domain inputs, synthetic perturbations, and black-box stress tests.
+Newer setups (e.g., MedQA-Adversarial, RAG robustness tests) probe noisy retrieval and unfamiliar conditions, but clinical realism and standardization remain limited.
+Limitations
+Lack of standardized, clinically grounded stress tests; overreliance on synthetic or narrow tasks; and brittle/costly mitigations (e.g., adversarial training, RAG) impede deployment.
+Multimodal and long-context robustness remain underexplored.
 3.5 Fairness and Bias
-Ensuring fairness in LLMs is crucial in healthcare, where biased models can result in unequal treatment outcomes.
-Research has highlighted biases in clinical data and practice related to race, gender, and disability.
+Findings in Fairness Bias across race, gender, and identity persists in medical LLMs.
+New benchmarks and mitigations help but are often narrow or misaligned with clinical reality.
+Progress requires intersectional audits, inclusive datasets, and collaboration with affected communities.
+Ensuring fairness is essential: biased outputs can exacerbate disparities in access, diagnosis, and treatment.
+Key areas of concern include demographic bias (e.g., race, gender, identity), automated detection of these biases, mitigation strategies based on model accessibility, and the need for ethical clarity and conceptual frameworks.
+Recent work spans benchmark creation, debiasing techniques, prompt interventions, and calls for more transparent fairness evaluations.
+Bias identification remains a foundational step.
+Studies show that LLMs can replicate and even amplify racial, gender, and identity-based biases.
 For example, Omiye et al.
-(2023) examine the potential for harmful or inaccurate race-based content in LLMs, while Zack et al.
-(2024) discusse how language models encode societal biases that can affect healthcare outcomes.
-These studies underscore the need for fairness in LLM development to ensure equitable healthcare delivery.
-Efforts to address bias focus on both detection and mitigation.
+(2023), Zack et al.
+(2024), and Kim et al.
+(2023) highlight persistent demographic biases in medical responses.
+Zhao et al.
+(2024b) find that diagnostic recommendations vary unfairly by demographic group, while Xie et al.
+(2024) reveal systematic inequities in outputs concerning race and LGBTQ+ identities.
+Patel et al.
+(2024) further demonstrate that LLMs can reinforce social and gender-based stereotypes in sensitive areas such as sexual and reproductive health, underscoring the risks in context-dependent medical interactions.
+Detection and benchmarking tools help quantify and monitor these disparities.
 Swaminathan et al.
-(2024) offers an automated method for detecting race-based medicine stereotypes, while BiasMedQA Schmidgall et al.
-(2024) benchmarks cognitive biases in medical tasks across multiple models, revealing varying bias resilience.
-Mitigation strategies, such as bias education, and one-shot and few-shot bias demonstrations, are proposed to reduce but not fully eliminate bias.
-Pfohl et al.
-(2024) introduce frameworks for assessing health equity-related harms in LLMs, including EquityMedQA, a dataset for equity-focused testing.
-Additionally, Wei et al.
-(2024) distinguishe between intrinsic fairness, rooted in model training, and behavioral fairness, which relates to model operation in real-world applications, advocating for both to ensure equitable outcomes.
-Bias is also explored in closed-source models.
-Zack et al.
-(2024) evaluate racial and gender biases in clinical scenarios, finding that models often amplify societal biases.
-Similarly, Adam et al.
-(2022) show that biased AI recommendations can affect emergency decisions, while Yang et al.
-(2024c) identify healthcare disparities in model predictions based on patient demographics.
-For open-source LLMs, techniques such as reinforcement learning with clinician feedback Zack et al.
-(2024) and data augmentation during pre-training Parray et al.
-(2023) enhance training data quality to reduce bias.
-In contrast, for closed-source models, where internal representations are inaccessible, strategies like instruction fine-tuning Singhal et al.
-(2023) and prompt engineering Schmidgall et al.
-(2024) are employed to improve fairness in outputs.
-Further, Zhang et al.
-(2020) investigate how LLM embeddings can encode biases, particularly in clinical tasks, and applies adversarial debiasing to mitigate disparities.
-Lin and Ng (2023) identify cognitive biases in BERT, while Ke et al.
-(2024) use a multi-agent framework to explore how LLMs can mitigate biases in clinical decision-making.
-Additionally, Zhu et al.
-(2023) introduce CI4MRC, a method that addresses name-related bias in Machine Reading Comprehension (MRC) tasks by applying a causal interventional paradigm.
-In addressing bias, solutions vary depending on whether the bias is at the individual level (e.g., name-related information) or dataset level (e.g., biased racial distributions).
-Open-sourced LLMs benefit from direct accessibility, enabling robust interventions to mitigate bias.
-In contrast, closed-sourced models, due to their inaccessibility, rely on methods such as instruction fine-tuning and external post-processing tools to refine model outputs and reduce bias.
+(2024) propose tools for identifying race-based stereotypes in medical Q&A.
+Benchmarks such as BiasMedQA Schmidgall et al.
+(2024), EquityMedQA Pfohl et al.
+(2024), and FairMedFM Jin et al.
+(2024) offer frameworks for testing model behavior across diverse patient profiles and clinical contexts.
+Mitigation strategies differ by model accessibility.
+For open-source models, techniques like adversarial debiasing Zhang et al.
+(2020), causal intervention (CI4MRC) Zhu et al.
+(2023), multi-agent collaboration Ke et al.
+(2024), and attribute-neutral modeling Hu et al.
+(2024) are applied to reduce bias.
+Data augmentation Parray et al.
+(2023) and bias-aware embedding assessments Lin and Ng (2023) provide further tools to enhance fairness in pretraining and inference.
+Closed-source models present unique challenges due to limited transparency.
+In these cases, fairness is addressed via instruction fine-tuning Singhal et al.
+(2023), external prompt engineering Schmidgall et al.
+(2024), or bounded personalization strategies Kirk et al.
+(2024), though these are less interpretable and harder to audit.
+Ethical and conceptual considerations also play a role.
+Wei et al.
+(2024) call for distinguishing between intrinsic and behavioral fairness, while Zhui et al.
+(2024) and Cotta and Maddison (2024) promote fairness through education and prompt design.
+Finally, Adam et al.
+(2022) and Yang et al.
+(2024b) warn that unchecked bias can distort care decisions and patient trust, emphasizing the stakes of fairness in real-world applications.
+Evaluations
+Assess disparities across subgroups using quantitative metrics (performance gaps, stereotyping scores) and qualitative audits; new benchmarks (FairMedFM, EquityMedQA) target equity in clinical recommendations, though intersectional analyses remain limited.
+Limitations
+Many evaluations overlook intersectional/institutional biases, adapt generic NLP methods without clinical causal context, and lack longitudinal assessment.
+Integration with other trust dimensions (robustness, privacy) is limited.
 3.6 Explanability
-The lack of explainability in LLMs poses a significant barrier to building trust with clinical practitioners, thereby restricting their adoption in real-world healthcare systems.
-To address this challenge, efforts in clinical practice and research have prioritized improving the transparency of LLMs by developing more effective explanation mechanisms and incorporating human oversight.
-For instance, Shariatmadari et al.
-(2024) enhance biomedical applications by integrating knowledge graphs with language models and visualizing attention probabilities to provide clear and interpretable explanations for model predictions.
-Elsborg and Salvatore (2023) employ local explanation models to generate intuitive, case-specific insights, further advancing the explainability of LLMs and fostering trust in their medical applications.
-Another critical area of research focuses on medical imaging explainability.
+Findings in Explainability Despite advances in rationales and attention maps, most tools lack clinical relevance and faithfulness.
+Methods often misalign with clinician reasoning; progress needs domain-specific frameworks plus causal/counterfactual explanations.
+Lack of explainability limits clinical trust.
+Recent research explores both intrinsic (model-integrated) and post-hoc (output-interpretation) techniques to make LLM reasoning more interpretable.
+These methods span a wide range of modalities, including text, graphs, tables, and images, and often incorporate domain-specific knowledge or human-centered reasoning to bridge model outputs and clinical expectations.
+Intrinsic explainability methods enhance transparency by aligning model attention with medical knowledge.
+For example, Shariatmadari et al.
+(2024) integrate knowledge graphs with attention visualization, while Wang et al.
+(2022) use Wikipedia-derived medical concepts to guide attention for code prediction, resulting in more concept-consistent outputs.
+Similarly, structure-to-text models like Pathway2Text Yang et al.
+(2022) convert biomedical graphs into interpretable narratives, supporting a more intuitive understanding of complex structured inputs.
+Post-hoc strategies focus on generating faithful rationales and justifications.
+FaReBio Fang et al.
+(2024) highlights how summarization faithfulness suffers with increased abstractiveness and introduces a benchmark to evaluate reasoning fidelity.
+In the molecular domain, LLM-GCE He et al.
+(2024) generates counterfactuals for Graph Neural Networks (GNNs) using dynamic feedback to ensure chemically valid, interpretable explanations.
+Several methods target zero-shot interpretability without task-specific fine-tuning.
+RAG-IM Mahbub et al.
+(2024) enables table-based clinical predictions with natural language justifications, while Liu et al.
+(2023a) embed ChatGPT into a diagnostic workflow with integrated interpretability components.
+Retrieval-based systems such as Retrieval + KG Ji et al.
+(2024) and DDCoT Zheng et al.
+(2023) further enhance reasoning by chaining knowledge-grounded prompts across modalities.
+Explainability in imaging and multimodal contexts is also gaining traction.
+MedThink Gai et al.
+(2025) fuses visual and textual inputs to improve multimodal reasoning, and MedExQA Kim et al.
+(2024) supplies detailed rationales for visual question answering.
 Ghosh et al.
-(2023) propose a method that iteratively decomposes a black-box (BB) model into interpretable expert models and a residual network, where expert models specialize in specific data subsets and explain their reasoning using First-Order Logic (FOL).
-Efforts to improve reasoning and multimodal integration in LLMs have also gained momentum.
-MedViLaM Xu et al.
-(2024) and MedThink Gai et al.
-(2024) focus on improving medical understanding and reasoning by integrating both textual and visual data for complex medical tasks such as question answering and medical image classification, aiming to provide a more holistic view for decision-making.
-In contrast, MedExQA Kim et al.
-(2024) prioritizes explainability in medical QA systems by providing multiple explanations for its responses.
-Causal Graphs Meet Thoughts Luo et al.
-(2025) and Retrieval and Reasoning on KGs Ji et al.
-(2024) explore how to enhance complex reasoning in LLMs through knowledge graphs (KGs), helping these models retrieve relevant information from structured sources to answer questions more effectively and with better justification.
-For more specialized applications, TOSRR Liu et al.
-(2025) introduce tree-organized self-reflective retrieval, which aims to improve performance in the niche area of Traditional Chinese Medicine (TCM), combining LLMs with self-reflection techniques for more accurate and contextually relevant responses.
-Studies like DDCoT Zheng et al.
-(2023) and Layered Chain-of-Thought Prompting Sanwal (2025) develop methods for multimodal reasoning across multiple agents, with DDCoT focusing on chain-of-thought prompting for structured reasoning across tasks.
-Lastly, frameworks like A ChatGPT Aided Explainable Framework Liu et al.
-(2023a) aim for zero-shot diagnosis and interpretation, ensuring more accessible AI support in real-world medical settings.
-The trend towards explainability enhancement encompasses both intrinsic and post-hoc approaches.
-Intrinsic methods, such as knowledge graphs and decomposed expert models, integrate explainability directly into the models, offering natural transparency.
-Post-hoc methods, including local explanations and textual justifications, elucidate model decisions after predictions, improving user comprehension and trust.
+(2023) decomposes black-box decisions into expert modules with first-order logic (FOL) reasoning.
+Evaluations
+Assessments use attribution heatmaps (e.g., LIME/SHAP), human-in-the-loop ratings, and contrastive/instruction-following tests focused on clarity, factual alignment, and clinical usefulness; standardized healthcare benchmarks remain scarce.
+Limitations
+Attribution tools have uncertain clinical validity, few studies show improved clinician decisions, and claims are rarely compared across models.
+Many methods don’t scale to large, multimodal, instruction-following LLMs, leaving faithfulness and practicality unresolved.
+3.7 Cost and Efficiency Considerations
+While large language models offer transformative potential for healthcare, their real-world deployment faces substantial cost and efficiency constraints.
+Large medical LLMs (e.g., GPT-4, Med-PaLM 2) are expensive to train, fine-tune, and operate, requiring significant compute, memory, and HIPAA-compliant infrastructure; latency and resource demands further hinder use in low-resource settings.
+These constraints restrict access to well-funded institutions and slow real-world adoption.
+Smaller open-source models lower inference costs and enable local/edge deployment but often reduce performance—especially in truthfulness, safety, and robustness.
+Scaling AI in healthcare therefore requires balancing trustworthiness with computational efficiency.
 4 Future Directions
-We have reviewed key trust challenges in LLMs and existing solutions. This section highlights current limitations and proposes future directions.
-While efforts to enhance LLM truthfulness in healthcare have advanced, gaps remain, including limited adaptability in hallucination mitigation and weak source attribution.
-Future research should prioritize model-agnostic post-processing, improved self-correction, enhanced uncertainty quantification, and real-time fact-checking via structured knowledge integration.
-Existing privacy safeguards, such as de-identification and federated learning, remain imperfect.
-Strengthening de-identification methods, reinforcing federated learning defenses, refining differential privacy, and exploring homomorphic encryption and real-time audits are crucial next steps.
-Safety remains a pressing concern due to adversarial attacks and excessive safety alignment-induced hallucinations.
-Future work should focus on robust validation mechanisms, refined safety alignment strategies, and comprehensive evaluation frameworks.
-Improving robustness remains critical, with adversarial testing and uncertainty quantification methods needing to better handle medical data complexities.
-Future research should focus on clinically relevant adversarial tests, enhancing uncertainty techniques, and improving instruction robustness.
-Despite progress in fairness, key gaps persist, including the need for comprehensive bias assessments and real-world testing of mitigation strategies.
-Standardizing fairness metrics, conducting real-world evaluations, and assessing long-term impacts on healthcare equity are critical for progress.
-Advancements in explainability have yet to bridge significant gaps.
-Future research should focus on integrating explainability into clinical workflows, developing interactive explanations, and improving multimodal integration to enhance transparency and trust.
-Recently, multi-agent frameworks like TriageAgent Lu et al.
-(2024) have been introduced to streamline complex clinical tasks through agent collaboration.
-By harnessing their capabilities, we can embed trustworthiness mitigation and evaluation within a multi-agent system, enabling proactive monitoring and intervention across key areas such as truthfulness, privacy, robustness, fairness, and explainability in healthcare LLMs.
-Addressing these gaps will ensure that LLMs can be effectively integrated into healthcare systems, improving their reliability, privacy, safety, fairness, and transparency.
+While core trust dimensions, truthfulness, privacy, robustness, fairness, explainability, and safety, have been the focus of recent work, emerging model paradigms such as multi-agent systems, multi-modal models, and small open-source LLMs introduce new trust challenges underexplored.
+Multi-Agent LLMs
+Multi-agent LLMs enable distributed reasoning through collaboration between specialized agents, offering improved robustness and self-correction.
+However, they also raise concerns around coordination, error propagation, and the interpretability of inter-agent communication.
+Trustworthy multi-agent systems will require protocols for communication, verification, and evaluation that ensure factual alignment and fairness.
+For example, Lu et al.
+(2024) introduce TriageAgent, a clinical multi-agent framework with role-specific LLMs for diagnosis and decision-making.
+While it shows benefits like structured collaboration and early stopping, it also reveals trust challenges, including inconsistent agent confidence, limited transparency, and error propagation—highlighting the need for stronger verification and alignment in high-stakes settings.
+Multimodal Foundation Models
+Multi-modal LLMs combine text, images, and structured data, better reflecting real-world clinical inputs but complicating trust evaluation.
+Challenges include cross-modal hallucination, misalignment, and reduced explainability.
+Addressing these issues will require modality-specific assessments, interpretable fusion strategies, and fairness testing across both textual and visual modalities.
+For example, Liu et al.
+(2024b) evaluate open-source multimodal LLMs for genomics and proteomics, highlighting issues with factual consistency and alignment across modalities—underscoring the importance of structured evaluation and interpretable model design in biomedical contexts.
+Small Open-Source LLMs
+Small open-source medical LLMs are gaining traction for their transparency, adaptability, and lower computational demands, making them attractive for deployment in resource-constrained or privacy-sensitive settings.
+However, their reduced capacity often leads to increased hallucinations, weaker safety alignment, and heightened privacy risks during fine-tuning on limited clinical data.
+Ensuring their trustworthiness requires lightweight hallucination mitigation, privacy-preserving training, and scalable evaluation pipelines.
+Despite their growing use, few studies directly examine these trust issues in small medical LLMs, as most existing research focuses on larger or general-purpose models, leaving a critical gap in the literature.
 5 Conclusion
-The integration of LLMs into healthcare holds great promise, but realizing their full potential requires addressing the critical challenges outlined in this survey.
-A key concern is truthfulness, as inaccuracies in medical LLMs pose serious risks to patient safety, making their detection and mitigation an ongoing research priority.
-Equally vital are privacy, safety, robustness, fairness, and explainability, ensuring responsible deployment in real-world clinical settings.
-While existing solutions show progress, much work remains to enhance the reliability, transparency, and ethical implications of LLMs in healthcare.
-Future research must refine these areas, balancing performance with trustworthiness while preventing LLM-based systems systems from worsening healthcare disparities.
-Comprehensive benchmarks, cross-disciplinary collaboration, and model accountability frameworks will be essential.
-Additionally, regulatory oversight and ethical guidelines must ensure LLM applications align with medical standards and patient rights.
-Ultimately, achieving safe and equitable AI-driven healthcare will require ongoing efforts to improve both technical capabilities and societal frameworks.
+As large language models continue to expand their role in healthcare, ensuring their trustworthiness remains a critical challenge.
+This survey reviewed six core dimensions—truthfulness, privacy, safety, robustness, fairness, and explainability—highlighting key methods, benchmarks, and limitations in current research.
+While recent advances have laid important groundwork, most existing solutions remain narrowly scoped and lack integration across dimensions, limiting their effectiveness in real-world clinical settings.
+Acknowledgment
+Our work is sponsored by NSF #2442253, NAIRR Pilot with PSC Neocortex and NCSA Delta, Commonwealth Cyber Initiative, Children’s National Hospital, Fralin Biomedical Research Institute (Virginia Tech), Sanghani Center for AI and Data Analytics (Virginia Tech), Virginia Tech Innovation Campus, and generous gifts from Nivida, Cisco, and the Amazon + Virginia Tech Center for Efficient and Robust Machine Learning.
 Limitations
@@
 Additionally, the survey focuses on current research without delving into the long-term societal and health impacts of LLM deployment, such as changes in doctor-patient relationships, patient trust, and healthcare workflows.
-Appendix A Comparison of Datasets
-We systematically collected and analyzed 41 datasets relevant to the study of trust in LLMs for healthcare.
+Appendix A Inclusion & Exclusion Criteria Details
+We conducted an extensive search to identify peer-reviewed papers that address the trustworthiness of LLMs in the healthcare domain.
+Our search strategy involved a wide range of keyword combinations related to LLMs and core trust dimensions, including: trustworthiness, truthfulness, privacy, safety, robustness, fairness, bias, and explainability.
+We targeted both domain-specific and general AI venues, focusing on recent publications from 2022 onward.
+Specifically, we searched across top-tier conferences and journals, including ACL, EMNLP, NAACL, ICML, NeurIPS, ICLR, KDD, AAAI, IJCAI, Nature, and Science, using platforms such as Google Scholar, Nature, and Science.
+A full list of keyword queries used in our search is provided below.
+These queries combined domain terms (medical, clinical) with trust-related dimensions, applied across both “large language models” and “foundation models.” Examples include:
+  • large language models, medical, explainability
+  • large language models, medical, explainable
+  • foundation model, medical, explainability
+  • large language models, clinical, explainability
+  • large language models, medical, truthfulness
+  • large language models, medical, trustworthiness
+  • foundation model, medical, trustworthiness
+  • large language models, clinical, truthfulness
+  • large language models, clinical, safety
+  • large language models, medical, safety
+  • foundation model, medical, safety
+  • large language models, clinical, fairness
+  • large language models, medical, fairness
+  • foundation model, medical, fairness
+  • large language models, clinical, robustness
+  • foundation model, medical, robustness
+  • large language models, medical, robustness
+  • large language models, clinical, privacy
+  • large language models, medical, privacy
+  • foundation model, medical, privacy
+  • large language models, clinical, ethics
+  • large language models, medical, ethics
+  • foundation model, medical, ethics
+In total, our initial search returned approximately 15,322 results, including duplicates and non-relevant papers. Our filtering process proceeded in three stages:
+  • Duplicate removal – approximately 11,172 papers eliminated.
+  • Relevance screening – we excluded papers that: (a) did not focus on trustworthiness aspects (e.g., architecture design or multi-modal fusion techniques), (b) were not specific to the healthcare domain, or (c) were unpublished preprints (e.g., arXiv manuscripts).
+  • Final selection – we curated a final set of 62 papers that directly addressed trust-related challenges in healthcare LLMs, focusing on one or more of the following dimensions: truthfulness, privacy, safety, robustness, fairness, bias, and explainability.
+Appendix B Comparison of Datasets
+We systematically collected and analyzed 38 datasets relevant to the study of trust in LLMs for healthcare.
 Table 1 provides a comprehensive summary, highlighting key attributes such as data type, content, associated tasks, and the specific trustworthiness dimensions they address.
@@
 Datasets Data Type Content Task Dimensions
-MultiMedQA Combination of Public and Synthetic Data, Curated Domain-Specific Dataset A benchmark combining six existing medical questions answering datasets spanning professional medicine, research and consumer queries and a new dataset of medical questions searched online, HealthSearchQA.
-Tasks including Medical Question Answering, Clinical Reasoning, Evidence-Based Medicine, Multilingual and Multimodal Support, Bias and Safety Analysis Fairness and Bias
-BiasMedQA Curated Domain-Specific Datasets 1273 USMLE questions Replicate common clinically relevant cognitive biases Fairness and Bias
-NEJM Healer Real Data and Curated Domain-Specific Dataset, Consists of Clinical Cases, Diagnostic Pathways, Educational Materials, Interactive Learning Modules Tasks including Diagnostic Skill Development, Medical Education, Simulated Decision-Making, Feedback and Improvement Fairness and Bias
-EquityMedQA Curated domain-specific datasets and synthetic data Cover a wide range of medical topics to surface biases that could harm health equity, including implicit and explicit adversarial questions addressing biases like stereotypes, lack of structural explanations, and withholding information.
-Evaluate the performance of LLMs in generating unbiased, equitable medical responses.
-Fairness and Bias
-”bias” medical dataset curated, domain-specific dataset comprising real-world healthcare data from a large U.S.
-health system Includes patient demographics, medical histories, and healthcare utilization records.
-Evaluate and identify racial biases in algorithms used for healthcare management.
+MultiMedQA Combination of Public and Synthetic Data, Curated Domain-Specific Dataset 208,000 entries.
+A benchmark combining six existing medical questions answering datasets spanning professional medicine, research and consumer queries and a new dataset of medical questions searched online, HealthSearchQA.
+(Med-QA) Tasks including Medical Question Answering, Clinical Reasoning, Evidence-Based Medicine, Multilingual and Multimodal Support, Bias and Safety Analysis Fairness and Bias
+BiasMedQA Curated Domain-Specific Datasets 1273 USMLE questions (Med-QA) Replicate common clinically relevant cognitive biases Fairness and Bias
+EquityMedQA Curated domain-specific datasets and synthetic data 4,619 examples.
+Cover a wide range of medical topics to surface biases that could harm health equity, including implicit and explicit adversarial questions addressing biases like stereotypes, lack of structural explanations, and withholding information.
+(Med-QA) Evaluate the performance of LLMs in generating unbiased, equitable medical responses.
 Fairness and Bias
@@
 Each question is paired with a segment of text from the corresponding article, serving as the answer.
-To develop models that can read a passage and answer questions about it, assessing the model’s ability to understand and extract information from the text.
+(Med-QA)To develop models that can read a passage and answer questions about it, assessing the model’s ability to understand and extract information from the text.
 Fairness and Bias
-MIMIC Real Data and Curated Domain-Specific Datasets Consists of electronic health records include patient Demographics, Clinical Data, Medical Notes, Treatment Records, Time-series Data Various medical and machine-learning tasks, including Clinical Decision Support, Disease Modeling, Natural Language Processing, Time-series Analysis, Education and Research Fairness and Bias, Explainability
-MedQA Curated Domain-Specific Datasets A benchmark that includes questions drawn from the United States Medical License Exam (USMLE).
-Exam the physicians to test their ability to make clinical decisions Fairness and Bias, Robustness, Explainability
+MIMIC- III Public text corpora, real-world data De-identified health-related data from over 40,000 critical care patients, including demographics, vital signs, laboratory tests, medications, and caregiver notes.
+(Med-IE) Epidemiological studies, clinical decision-rule improvement, machine learning in healthcare.
+Fairness and Bias, Explainability, Robustness
+MedQA Curated Domain-Specific Datasets 194,000 multiple-choice medical exam questions.
+A benchmark that includes questions drawn from the United States Medical License Exam (USMLE).
+(Med-QA) Exam the physicians to test their ability to make clinical decisions Fairness and Bias, Robustness, Explainability, Truthfulness, Privacy
 PMC-Patients Curated dataset derived from public text corpora.
-Contains 167,000 patient summaries extracted from 141,000 PMC articles Designed to benchmark ReCDS systems through two primary tasks: Patient-to-Article Retrieval (PAR), Patient-to-Patient Retrieval (PPR) Robustness
-MIMIC- III Public text corpora, real-world data De-identified health-related data from over 40,000 critical care patients, including demographics, vital signs, laboratory tests, medications, and caregiver notes.
-Epidemiological studies, clinical decision-rule improvement, machine learning in healthcare.
-Robustness
-Datasets Data Type Content Task Dimensions
+Contains 167,000 patient summaries extracted from 141,000 PMC articles (Med-IE) Designed to benchmark ReCDS systems through two primary tasks: Patient-to-Article Retrieval (PAR), Patient-to-Patient Retrieval (PPR) Robustness
 MedSafetyBench Curated domain-specific dataset and synthetic (generated using GPT-4, Llama-2-7b-chat, and adversarial techniques).
@@
 The dataset is structured based on the Principles of Medical Ethics from the American Medical Association (AMA).
-Assess the medical safety of LLMs by testing whether they refuse to comply with harmful medical requests.
+(Med-Gen) Assess the medical safety of LLMs by testing whether they refuse to comply with harmful medical requests.
 Fine-tune LLMs using medical safety demonstrations to enhance their alignment with ethical medical guidelines.
 Safety
-UNIWIZ Synthetic and curated data, including: 17,638 quality-controlled conversations, and 10,000 augmented preference data Features conversations that integrate safety and knowledge alignment.
+Datasets Data Type Content Task Dimensions
+UNIWIZ Synthetic and curated data, including: 17,638 quality-controlled conversations, and 10,000 augmented preference data 17,638 conversations and 10,000 augmented preference data.
+Features conversations that integrate safety and knowledge alignment.
 A "safety-priming" method was employed to generate synthetic safety data, and factual information was injected into conversations by retrieving content from curated sources.
-Fine-tune large language models to enhance their performance in generating safe and knowledge-grounded conversations.
+(Med-Gen) Fine-tune large language models to enhance their performance in generating safe and knowledge-grounded conversations.
 Safety
 SciFact Curated Domain-Specific Dataset.
+2,011 claims.
 Includes claims and corresponding evidence abstracts, each annotated with labels indicating whether the claim is supported or refuted, along with rationales justifying the decision.
-To verify the veracity of scientific claims by identifying supporting or refuting evidence within abstracts and providing justifications for these decisions.
+(Med-Gen) To verify the veracity of scientific claims by identifying supporting or refuting evidence within abstracts and providing justifications for these decisions.
 Truthfulness
 PubHealthTab Curated Domain-Specific Dataset Contains 1,942 real-world public health claims, each paired with evidence tables extracted from over 300 websites.
-Facilitates evidence-based fact-checking by providing claims and corresponding evidence tables for verification.
+(Med-Gen) Facilitates evidence-based fact-checking by providing claims and corresponding evidence tables for verification.
 Truthfulness
 LAMA Curated Domain-Specific Dataset.
+24,223 entries of knowledge sources.
 Comprises a set of knowledge sources, each containing a collection of facts.
-To probe pretrained language models to determine the extent of their factual and commonsense knowledge.
+(Med-Gen) To probe pretrained language models to determine the extent of their factual and commonsense knowledge.
 Truthfulness
@@
 The questions are sourced from trivia websites, and the answers are derived from the corresponding documents.
-Training and evaluating models on reading comprehension, specifically focusing on the ability to extract and reason over information from provided documents to answer questions.
+(Med-QA) Training and evaluating models on reading comprehension, specifically focusing on the ability to extract and reason over information from provided documents to answer questions.
 Truthfulness
-Natural Questions (NQ) Real data consists of real anonymized queries from Google’s search engine users, paired with answers derived from entire Wikipedia articles.
-To develop and evaluate question-answering systems that can read and comprehend entire Wikipedia articles to find answers to user queries.
+Natural Questions (NQ) Real data  99.80 GB, with downloaded files accounting for 45.07 GB and the generated dataset occupying 54.73 GB.
+consists of real anonymized queries from Google’s search engine users, paired with answers derived from entire Wikipedia articles.
+(Med-QA) To develop and evaluate question-answering systems that can read and comprehend entire Wikipedia articles to find answers to user queries.
 Truthfulness
@@
 consists of 14,000 QA pairs, each associated with fine-grained Wikidata entity IDs, Wikipedia page views, and relationship type information.
-Designed for open-domain question answering tasks, focusing on evaluating the effectiveness of language models in retrieving and utilizing factual knowledge.
+(Med-QA) Designed for open-domain question answering tasks, focusing on evaluating the effectiveness of language models in retrieving and utilizing factual knowledge.
 Truthfulness
@@
 These claims are categorized as supported, refuted, or not verifiable.
-Fact extraction and verification, where models are trained to determine the veracity of claims based on provided evidence.
-Truthfulness
-HEALTHVER Curated Domain-Specific Dataset.
-Contains 14,330 evidence-claim pairs (SUPPORTS, REFUTES, NEUTRAL) on health claims, mainly COVID-19, verified against scientific articles.
-Used to train and evaluate models in verifying health claims by classifying them based on scientific evidence.
+(Med-Gen) Fact extraction and verification, where models are trained to determine the veracity of claims based on provided evidence.
 Truthfulness
 Datasets Data Type Content Task Dimensions
-Med-HALT Synthetic and Real Data, Curated Domain-Specific Dataset, and Public Dataset Consist of Reasoning-Based Assessments, Memory-Based Assessments, Medical Scenarios, Evaluation Metrics Tasks including Evaluation of Hallucination in Medical AI, Reliability Benchmarking, Error Analysis, Mitigation Development Truthfulness
-MedICaT Public Text Corpora And Real Data (curated from publicly available biomedical literature) Contains medical images (e.g., radiographs, charts, and diagrams) paired with captions extracted from biomedical literature.
+HEALTHVER Curated Domain-Specific Dataset.
+contains 14,330 evidence-claim pairs labeled as SUPPORTS, REFUTES, or NEUTRAL, derived from real-world health claims, mainly about COVID-19, verified against scientific articles.
+(Med-Gen) Training and evaluating models on the task of verifying the truthfulness of health-related claims by assessing their alignment with scientific evidence.
+This involves classifying claims as supported, refuted, or neutral based on the provided evidence.
+Truthfulness
+Med-HALT Synthetic and Real Data, Curated Domain-Specific Dataset, and Public Dataset 59,254 entries.
+Consist of Reasoning-Based Assessments, Memory-Based Assessments, Medical Scenarios, Evaluation Metrics (Med-Gen) Tasks including Evaluation of Hallucination in Medical AI, Reliability Benchmarking, Error Analysis, Mitigation Development Truthfulness
+MedICaT Public Text Corpora And Real Data (curated from publicly available biomedical literature) 217,060 figures extracted from 131,410 open-access papers.
+Contains medical images (e.g., radiographs, charts, and diagrams) paired with captions extracted from biomedical literature.
 Also, includes metadata about the source and context of the images.
-Task including Medical Image Captioning, Text-Image Retrieval, Medical Reasoning Truthfulness
-Med-HallMark Curated Domain-Specific Dataset, , Synthetic and Real Data (includes a mix of real-world medical data and synthetically generated hallucination scenarios) Diverse medical multimodal data, including text, images, and paired annotations.
-Hierarchically categorized hallucination data, addressing both structural (e.g., object-level) and contextual (e.g., domain knowledge) hallucinations.
-Task including Hallucination Detection, Hallucination Evaluation, Mitigation Analysis Truthfulness
+(Med-Gen) Task including Medical Image Captioning, Text-Image Retrieval, Medical Reasoning Truthfulness
 BioASQ Curated Domain-Specific Dataset; Real Data.
+3,743 training questions and 500 test questions.
 The dataset comprises English-language biomedical questions, each accompanied by reference answers and related materials.
 These questions are designed to reflect real information needs of biomedical experts, making the dataset both realistic and challenging.
-The primary task is Biomedical Question Answering (QA), which involves systems providing accurate answers to questions based on biomedical data.
+(Med-QA) The primary task is Biomedical Question Answering (QA), which involves systems providing accurate answers to questions based on biomedical data.
 The dataset supports various QA tasks, including yes/no, factoid, list, and summary questions.
@@
 The dataset was generated using GPT-4 and includes biographies of 21 individuals randomly sampled from the WikiBio dataset.
-Evaluating the factual accuracy of language models, particularly in the context of biography generation.
+(Med-Gen) Evaluating the factual accuracy of language models, particularly in the context of biography generation.
 It serves as a benchmark for detecting hallucinations and assessing the factual consistency of generated text.
@@
 Consists of over 1,000 question-answer pairs derived from PubMed abstracts, focusing on various biomedical topics.
-Evaluates the ability of models to comprehend and extract information from biomedical texts to answer specific questions.
+(Med-QA) Evaluates the ability of models to comprehend and extract information from biomedical texts to answer specific questions.
 Truthfulness
@@
 The dataset encompasses 37 question types, such as Treatment, Diagnosis, and Side Effects, associated with diseases, drugs, and other medical entities like tests.
-Designed for medical question answering, the dataset aids in developing and evaluating systems that can understand and respond to medical inquiries.
+(Med-QA) Designed for medical question answering, the dataset aids in developing and evaluating systems that can understand and respond to medical inquiries.
 Truthfulness
-LiveMedQA2017 Curated Domain-Specific Dataset Consists of 634 question-answer pairs corresponding to National Library of Medicine (NLM) questions Medical question answering, focusing on consumer health questions received by the U.S.
+LiveMedQA2017 Curated Domain-Specific Dataset Consists of 634 question-answer pairs corresponding to National Library of Medicine (NLM) questions (Med-QA) Medical question answering, focusing on consumer health questions received by the U.S.
 National Library of Medicine.
 Truthfulness
+Datasets Data Type Content Task Dimensions
 MASH-QA Curated Domain-Specific Dataset.
 Approximately 25,000 question-answer pairs sourced from WebMD, covering a wide range of healthcare topics.
-Designed for multiple-answer span extraction in healthcare question answering.
+(Med-QA) Designed for multiple-answer span extraction in healthcare question answering.
 Truthfulness
@@
 The dataset encompasses 57 databases across 34 diverse domains, each associated with specific security conditions.
-Evaluate and analyze data leakage risks in LLMs, particularly concerning SQL query generation and execution.
+(Med-IE) Evaluate and analyze data leakage risks in LLMs, particularly concerning SQL query generation and execution.
 Privacy
+Medical Meadow curated domain-specific dataset It comprises approximately 1.5 million data points across various tasks, including question-answer pairs generated from openly available medical data using models like OpenAI’s (Med-Gen) Designed to enhance large language models (LLMs) for medical applications Privacy
+Electronic Health Records (EHR) at (KHCC) Private dataset gpt-3.5-turbo (Med-IE) Clinical research, outcome analysis. Privacy
+MedVQA Curated domain-specific dataset 794 image-question-answer triplets.
+A collection of medical visual question answering pairs, designed to train and evaluate models that interpret medical images and answer related questions.
+(Med-QA) Visual question answering, medical image understanding.
+Explainability
+MedExQA Curated domain-specific dataset 965 multiple-choice medical questions.
+A dataset focused on medical examination questions and answers, intended to aid in the development of AI models for medical exam preparation and assessment.
+(Med-QA) Question answering, educational assessment.
+Explainability
+MedMCQA Curated domain-specific dataset 194,000 multiple-choice questions from AIIMS and NEET PG entrance exams, covering 2,400 healthcare topics across 21 medical subjects.
+A multiple-choice question-answering dataset in the medical domain, aimed at training models to handle medical examinations and practice questions.
+(Med-QA) Multiple-choice question answering, medical education.
+Explainability
+TCM Medical Licensing Examination(MLE) Curated domain-specific dataset 600 multiple-choice questions.
+A dataset comprising questions and answers from Traditional Chinese Medicine licensing examinations.
+(Med-QA) Educational assessment, question answering.
+Explainability
 Datasets Data Type Content Task Dimensions
-Medical Meadow curated domain-specific dataset It comprises approximately 1.5 million data points across various tasks, including question-answer pairs generated from openly available medical data using models like OpenAI’s Designed to enhance large language models (LLMs) for medical applications Privacy
-Electronic Health Records (EHR) at (KHCC) Private dataset gpt-3.5-turbo Clinical research, outcome analysis. Privacy
-MedVQA Curated domain-specific dataset A collection of medical visual question answering pairs, designed to train and evaluate models that interpret medical images and answer related questions.
-Visual question answering, medical image understanding.
+Pneumonia Dataset Curated domain-specific dataset 5,863 images.
+Medical images (such as chest X-rays) labeled for the presence or absence of pneumonia, used for training diagnostic models.
+(Med-IE) Image classification, disease detection.
 Explainability
-MedExQA Curated domain-specific dataset A dataset focused on medical examination questions and answers, intended to aid in the development of AI models for medical exam preparation and assessment.
-Question answering, educational assessment.
+Montgomery Dataset Curated domain-specific dataset X-ray Set comprises 138 posterior-anterior chest X-ray images, with 80 normal and 58 abnormal cases indicative of tuberculosis.
+Chest X-ray images with manual segmentations of the lung fields, useful for pulmonary research.
+(Med-IE) Image segmentation, tuberculosis detection.
 Explainability
-MedMCQA Curated domain-specific dataset A multiple-choice question-answering dataset in the medical domain, aimed at training models to handle medical examinations and practice questions.
-Multiple-choice question answering, medical education.
+Shenzhen Dataset Curated domain-specific dataset Chest X-ray dataset comprises 662 frontal chest X-rays, including 326 normal cases and 336 cases with manifestations of tuberculosis.
+Chest X-ray images collected in Shenzhen, China, with annotations for tuberculosis manifestations.
+(Med-IE) Disease classification, image analysis.
 Explainability
-TCM Medical Licensing Examination(MLE) Curated domain-specific dataset A dataset comprising questions and answers from Traditional Chinese Medicine licensing examinations.
-Educational assessment, question answering.
+IDRID Dataset Curated domain-specific dataset 1,113 images.
+Retinal images with annotations for diabetic retinopathy lesions, intended for retinal image analysis.
+(Med-IE) Image segmentation, disease grading.
 Explainability
-Pneumonia Dataset Curated domain-specific dataset Medical images (such as chest X-rays) labeled for the presence or absence of pneumonia, used for training diagnostic models.
-Image classification, disease detection.
-Explainability
-Montgomery Dataset Curated domain-specific dataset Chest X-ray images with manual segmentations of the lung fields, useful for pulmonary research.
-Image segmentation, tuberculosis detection.
-Explainability
-Shenzhen Dataset Curated domain-specific dataset Chest X-ray images collected in Shenzhen, China, with annotations for tuberculosis manifestations.
-Disease classification, image analysis.
-Explainability
-IDRID Dataset Curated domain-specific dataset Retinal images with annotations for diabetic retinopathy lesions, intended for retinal image analysis. Image segmentation, disease grading. Explainability
-BrainTumor Dataset Curated domain-specific dataset MRI images of brain tumors with corresponding labels, used for developing diagnostic and segmentation models.
-Tumor detection, image segmentation.
-Explainability
+MIMIC IV Curated Real-World Clinical Dataset Over 300,000 hospital admissions from Beth Israel Deaconess Medical Center covering de-identified EHR data including demographics, vital signs, medications, diagnoses, and clinical notes (Med-IE / Med-QA / Med-Gen) Used for tasks such as medical code prediction, patient outcome forecasting, clinical summarization, and question answering Explainability
 Table 1: This table provides a structured comparison of datasets used in studies on trust in LLMs for healthcare.
@@
 This comparison highlights how each dataset contributes to the development of trustworthy LLMs in medical AI.
-Appendix B Comparison of Models
-We systematically gathered and analyzed 78 models relevant to studies on trust in LLMs for healthcare.
-Table 2 provides a comprehensive summary of the LLMs evaluated in these studies, detailing key aspects such as model name, release year, and the institution responsible for its development.
+Appendix C Comparison of Models
+We systematically gathered and analyzed 81 models relevant to studies on trust in LLMs for healthcare.
+Table 2 provides a comprehensive summary of the LLMs evaluated in these studies, detailing key aspects such as model name, release year, openness, architecture, and the institution responsible for its development.
 Additionally, it specifies the primary task each model is designed for, including medical question-answering, clinical decision support, and biomedical text summarization.
@@
 This structured overview offers valuable insights into how different LLMs are designed and evaluated to enhance trust in healthcare AI applications.
-Models Release Year Institution Primary Task Dimensions
-SciBERT 2019 Allen Institute for AI Pre-trained language model specialized for scientific text, particularly biomedical and computer science literature. Fairness and Bias
-PaLM-2 2023 Google Multilingual language understanding and generation, with a focus on reasoning and coding tasks. Fairness and Bias
-Mixtral-8x70B 2023 Mistral AI Ensemble of language models aimed at improving performance across diverse language tasks. Fairness and Bias, Safety
-Med-PaLM 2023 Google Health Specializing in healthcare-related question answering, clinical diagnosis support, and medical literature interpretation. Fairness and Bias
-Med-PaLM 2 2024 Google Health Updated version of Med-PaLM, further improving healthcare-related tasks with enhanced accuracy and reliability in medical information retrieval, clinical reasoning, and decision support.
+Models Release Year Institution Openness Architecture Primary Task Dimensions
+SciBERT 2019 Allen Institute for AI Open-source Encoder-only Pre-trained language model specialized for scientific text, particularly biomedical and computer science literature. Fairness and Bias
+PaLM-2 2023 Google Closed-source Decoder-only Multilingual language understanding and generation, with a focus on reasoning and coding tasks. Fairness and Bias
+Mixtral-8x70B 2023 Mistral AI Open-source Decoder-only Ensemble of language models aimed at improving performance across diverse language tasks. Fairness and Bias, Safety
+Med-PaLM 2023 Google Health Closed-source Decoder-only Specializing in healthcare-related question answering, clinical diagnosis support, and medical literature interpretation. Fairness and Bias
+Med-PaLM 2 2024 Google Health Closed-source Encoder-decoder Updated version of Med-PaLM, further improving healthcare-related tasks with enhanced accuracy and reliability in medical information retrieval, clinical reasoning, and decision support.
 Fairness and Bias
-Llama-13B 2023 Meta Designed for natural language understanding and generation tasks, such as text summarization, machine translation, and conversational AI. Fairness and Bias
-XLNet 2019 Google Research It is used for text classification, question answering, and language modeling tasks. Fairness and Bias
-DeBERTa 2020 Microsoft Research Improves BERT and RoBERTa by enhancing the attention mechanism.
+Llama-13B 2023 Meta Open-source Decoder-only Designed for natural language understanding and generation tasks, such as text summarization, machine translation, and conversational AI. Fairness and Bias
+XLNet 2019 Google Research Open-source Encoder-only It is used for text classification, question answering, and language modeling tasks. Fairness and Bias
+DeBERTa 2020 Microsoft Research Open-source Encoder-only Improves BERT and RoBERTa by enhancing the attention mechanism.
 It performs well in a variety of NLP tasks, such as sentence classification, question answering, and named entity recognition.
 Fairness and Bias
-Llama-7B 2023 Meta Focused on general-purpose natural language understanding and generation, with potential fine-tuning for specific domains like medicine, law, and technology.
+Llama-7B 2023 Meta Open-source Decoder-only Focused on general-purpose natural language understanding and generation, with potential fine-tuning for specific domains like medicine, law, and technology.
 Fairness and Bias, Truthfulness
-Llama 2 70Bchat 2023 Meta Platforms Open-source conversational AI model designed for dialogue and instruction-following tasks. Fairness and Bias, Truthfulness, Safety, Robustness,
-GPT-3.5 2022 OpenAI Enhanced language processing capabilities, building upon GPT-3. Fairness and Bias, Truthfulness, Safety, Robustness, Privacy
-GPT2 2019 OpenAI Text generation Fairness and Bias, Robustness
-PMC Llama 13B 2023 Allen Institute for AI Specialized in medical literature understanding and generation. Fairness and Bias, Robustness
-Models Release Year Institution Primary Task Dimensions
-GPT-4 2023 OpenAI Advanced language generation and understanding across various domains. Fairness and Bias, Safety, Robustness, Explainability, Privacy
-BERT 2018 Google AI Language Pre-trained Transformer model for a wide range of NLP tasks, such as text classification, NER, QA, etc. Fairness and Bias, Safety, Robustness, Truthfulness
-LLAMA 2 CHAT 2023 Meta AI Language modeling Robustness, Explainability
-MEDALPACA (7B) 2023 medalpaca Medical domain language model fine-tuned for question-answering and medical dialogue tasks. Robustness, Privacy
-CLINICAL CAMEL (13B) 2023 the AI and healthcare community Fine-tuned for clinical applications.
+Llama 2 70Bchat 2023 Meta Platforms Open-source Decoder-only Open-source conversational AI model designed for dialogue and instruction-following tasks.
+Fairness and Bias, Truthfulness, Safety, Robustness,
+GPT-3.5 2022 OpenAI Closed-source Decoder-only Enhanced language processing capabilities, building upon GPT-3. Fairness and Bias, Truthfulness, Safety, Robustness, Privacy
+GPT2 2019 OpenAI Open-source Decoder-only Text generation Fairness and Bias, Robustness
+Models Release Year Institution Openness Architecture Primary Task Dimensions
+PMC Llama 13B 2023 Allen Institute for AI Open-source Decoder-only Specialized in medical literature understanding and generation. Fairness and Bias, Robustness
+GPT-4 2023 OpenAI Closed-source Decoder-only Advanced language generation and understanding across various domains. Fairness and Bias, Safety, Robustness, Explainability, Privacy
+BERT 2018 Google AI Language Open-source Encoder-only Pre-trained Transformer model for a wide range of NLP tasks, such as text classification, NER, QA, etc.
+Fairness and Bias, Safety, Robustness, Truthfulness
+LLAMA 2 CHAT 2023 Meta AI Open-source Decoder-only Language modeling Robustness, Explainability
+MEDALPACA (7B) 2023 medalpaca Open-source Decoder-only Medical domain language model fine-tuned for question-answering and medical dialogue tasks. Robustness, Privacy
+CLINICAL CAMEL (13B) 2023 the AI and healthcare community Open-source Decoder-only Fine-tuned for clinical applications.
 It is designed to assist with tasks like medical text classification, clinical decision support, information extraction from medical records, and answering clinical questions.
 Robustness
-GPT-2 XL 2019 OpenAI Large-scale language model for text generation and understanding. Robustness
-T5-Large 2020 Google Research It treats all NLP tasks as text-to-text tasks, meaning both the input and output are in the form of text, and it’s used for tasks like translation, summarization, and question answering.
+GPT-2 XL 2019 OpenAI Open-source Decoder-only Large-scale language model for text generation and understanding. Robustness
+T5-Large 2020 Google Research Open-source Encoder-decoder It treats all NLP tasks as text-to-text tasks, meaning both the input and output are in the form of text, and it’s used for tasks like translation, summarization, and question answering.
 Robustness
-claude-3.5-sonnet 2024 Anthropic It is a variant of Claude, specialized in tasks such as conversational AI, creative writing, poetry generation, and other text-based applications. Robustness
-OpenBioLLM-70B 2024 OpenBioAI It is designed to handle tasks such as biological information extraction, gene sequence analysis, protein folding predictions, and other bioinformatics applications.
+claude-3.5-sonnet 2024 Anthropic Closed-source Decoder-only It is a variant of Claude, specialized in tasks such as conversational AI, creative writing, poetry generation, and other text-based applications.
 Robustness
-BioMistral-7B 2023 Mistral AI Focused on biomedical and healthcare-related text. Its tasks include medical question answering, clinical document analysis, and medical text summarization. Robustness
-Medllama3-v20 2024 MedAI Labs Designed to assist in healthcare tasks like clinical reasoning, medical question answering, and patient record analysis. Robustness
-ASCLEPIUS (7B) 2023 Asclepius AI Developed for clinical and medical applications, specializing in tasks like diagnosing medical conditions from symptoms, medical text summarization, and extracting structured information from clinical documents.
+OpenBioLLM-70B 2024 OpenBioAI Open-source Decoder-only It is designed to handle tasks such as biological information extraction, gene sequence analysis, protein folding predictions, and other bioinformatics applications.
+Robustness
+BioMistral-7B 2023 Mistral AI Open-source Decoder-only Focused on biomedical and healthcare-related text.
+Its tasks include medical question answering, clinical document analysis, and medical text summarization.
+Robustness
+Medllama3-v20 2024 MedAI Labs Open-source Decoder-only Designed to assist in healthcare tasks like clinical reasoning, medical question answering, and patient record analysis. Robustness
+Models Release Year Institution Openness Architecture Primary Task Dimensions
+ASCLEPIUS (7B) 2023 Asclepius AI Open-source Decoder-only Developed for clinical and medical applications, specializing in tasks like diagnosing medical conditions from symptoms, medical text summarization, and extracting structured information from clinical documents.
 Robustness, Explainability
-ALPACA (7B) 2023 Stanford University Fine-tuned version of the LLaMA model aimed at providing high-quality responses to questions, with an emphasis on maintaining ethical and accurate conversational capabilities in diverse domains.
+ALPACA (7B) 2023 Stanford University Open-source Decoder-only Fine-tuned version of the LLaMA model aimed at providing high-quality responses to questions, with an emphasis on maintaining ethical and accurate conversational capabilities in diverse domains.
 Robustness
-Models Release Year Institution Primary Task Dimensions
-Google’s Bard 2023 Google Conversational AI tool, focused on providing detailed, accurate, and creative responses to user queries.
+Google’s Bard 2023 Google Closed-source Encoder-decoder Conversational AI tool, focused on providing detailed, accurate, and creative responses to user queries.
 It can handle a variety of tasks, including web search, content generation, and complex QA.
 Robustness
-Text- Davinci-003 2022 OpenAI It is an advanced variant of GPT-3.
+Text- Davinci-003 2022 OpenAI Closed-source Decoder-only It is an advanced variant of GPT-3.
 It is designed for a wide range of natural language understanding and generation tasks, such as answering questions, summarizing text, creative writing, translation, and code generation.
 Robustness, Truthfulness
-LLaMa 2-7B 2023 Meta (formerly Facebook AI Research) Designed to be a general-purpose AI for a wide range of tasks such as text generation, question answering, and summarization, with specific fine-tuning for medical and technical domains.
+LLaMa 2-7B 2023 Meta (formerly Facebook AI Research) Open-source Decoder-only Designed to be a general-purpose AI for a wide range of tasks such as text generation, question answering, and summarization, with specific fine-tuning for medical and technical domains.
 Robustness, Truthfulness, Privacy
-ChatGPT 2022 OpenAI Conversational AI Robustness, Truthfulness, Explainability, Privacy
-Llama-3.1 2024 Meta AI Multilingual large language model designed for a variety of natural language processing tasks. Safety
-ClinicalCamel-70b 2023 the AI and healthcare community Medical language model designed for clinical research applications. Safety, Explainability
-Med42-70b 2023 M42 Health Clinical large language model providing high-quality answers to medical questions. Safety, Explainability
-GPT-4o 2024 OpenAI Multimodal large language model capable of processing and generating text, audio, and images in real time. Safety, Privacy, Explainability
-Mistral 2023 Mistral AI Language model optimized for code generation and reasoning tasks. Safety, Robustness, Explainability
-Meditron (7) (70b) 2023 École Polytechnique Fédérale de Lausanne (EPFL) Medical language model fine-tuned for clinical decision support and medical reasoning. Safety, Robustness, Explainability
-Claude-2.1 2023 Anthropic General-purpose language model for a wide range of natural language understanding and generation tasks. Safety, Robustness
-GPT-J 2021 EleutherAI Open-source language model for text generation and understanding. Safety, Robustness
-Vicuna 2023 UC Berkeley and Microsoft Research Conversational AI Safety, Robustness, Truthfulness
-Models Release Year Institution Primary Task Dimensions
-Medalpaca-13b 2023 medalpaca Medical domain language model fine-tuned for question-answering and medical dialogue tasks. Safety, Truthfulness, Privacy
-GPT-3 2020 OpenAI Natural language understanding and generation Truthfulness, Explainability
-ALBERT 2019 Google Research Lighter version of BERT that reduces parameters for efficiency while maintaining performance.
+ChatGPT 2022 OpenAI Closed-source Decoder-only Conversational AI Robustness, Truthfulness, Explainability, Privacy
+Llama-3.1 2024 Meta AI Open-source Decoder-only Multilingual large language model designed for a variety of natural language processing tasks. Safety, privacy
+ClinicalCamel-70b 2023 the AI and healthcare community Open-source Decoder-only Medical language model designed for clinical research applications. Safety, Explainability
+Med42-70b 2023 M42 Health Open-source Decoder-only Clinical large language model providing high-quality answers to medical questions. Safety, Explainability
+GPT-4o 2024 OpenAI Closed-source Decoder-only Multimodal large language model capable of processing and generating text, audio, and images in real time. Safety, Privacy, Explainability
+Mistral 2023 Mistral AI Open-source Decoder-only Language model optimized for code generation and reasoning tasks. Safety, Robustness, Explainability
+Meditron (7) (70b) 2023 École Polytechnique Fédérale de Lausanne (EPFL) Open-source Decoder-only Medical language model fine-tuned for clinical decision support and medical reasoning.
+Safety, Robustness, Explainability
+Models Release Year Institution Openness Architecture Primary Task Dimensions
+Claude-2.1 2023 Anthropic Closed-source Decoder-only General-purpose language model for a wide range of natural language understanding and generation tasks. Safety, Robustness
+GPT-J 2021 EleutherAI Open-source Decoder-only Open-source language model for text generation and understanding. Safety, Robustness
+Vicuna 2023 UC Berkeley and Microsoft Research Open-source Decoder-only Conversational AI Safety, Robustness, Truthfulness
+Medalpaca-13b 2023 medalpaca Open-source Decoder-only Medical domain language model fine-tuned for question-answering and medical dialogue tasks. Safety, Truthfulness, Privacy
+GPT-3 2020 OpenAI Closed-source Decoder-only Natural language understanding and generation Truthfulness, Explainability
+ALBERT 2019 Google Research Open-source Encoder-only Lighter version of BERT that reduces parameters for efficiency while maintaining performance.
 It excels in tasks such as text classification, named entity recognition, and question answering.
 Truthfulness
-RoBERTa 2019 Facebook AI Research Optimized variant of BERT that removes the Next Sentence Prediction task and trains with more data and for longer periods.
+RoBERTa 2019 Facebook AI Research Open-source Encoder-only Optimized variant of BERT that removes the Next Sentence Prediction task and trains with more data and for longer periods.
 It is used for tasks like question answering, sentiment analysis, and text classification.
 Truthfulness
-BlueBERT 2019 NIH and Stanford University BERT-based model pre-trained on clinical and biomedical text.
+BlueBERT 2019 NIH and Stanford University Open-source Encoder-only BERT-based model pre-trained on clinical and biomedical text.
 It is designed for healthcare-related tasks, including clinical text classification, named entity recognition, and medical question answering.
 Truthfulness
-ClinicalBERT 2019 University of Pennsylvania Variant of BERT fine-tuned on clinical texts, tailored for clinical NLP tasks like named entity recognition, clinical event extraction, and question answering in the medical domain.
+ClinicalBERT 2019 University of Pennsylvania Open-source Encoder-only Variant of BERT fine-tuned on clinical texts, tailored for clinical NLP tasks like named entity recognition, clinical event extraction, and question answering in the medical domain.
 Truthfulness
-TAPAS 2020 Google Research Designed for answering questions based on tabular data.
+TAPAS 2020 Google Research Open-source Encoder-only Designed for answering questions based on tabular data.
 It is used for tasks like extracting structured information from tables and processing queries in tabular datasets.
 Truthfulness
-LLaMA-2 13B 2023 Meta Advanced variant of Meta’s LLaMA series, designed for text generation, question answering, summarization, and other NLP tasks. Truthfulness, Explainability, Privacy
-MPT 2023 MosaicML General-purpose LLM for text generation, summarization, language understanding, and reasoning tasks.
+LLaMA-2 13B 2023 Meta Open-source Decoder-only Advanced variant of Meta’s LLaMA series, designed for text generation, question answering, summarization, and other NLP tasks.
+Truthfulness, Explainability, Privacy
+MPT 2023 MosaicML Open-source Decoder-only General-purpose LLM for text generation, summarization, language understanding, and reasoning tasks.
 Fine-tuned for downstream applications such as chatbot development, code generation, and other NLP tasks.
 Truthfulness
-BLIP2 2023 Salesforce Bootstrapping language-image pre-training, designed to bridge vision-language models with large language models for improved visual understanding and generation. Truthfulness
-InstructBLIP-7b/13b 2023 Salesforce Visual instruction-tuned versions of BLIP-2, utilizing Vicuna-7B and Vicuna-13B language models, respectively, to enhance vision-language understanding through instruction tuning.
+Models Release Year Institution Openness Architecture Primary Task Dimensions
+BLIP2 2023 Salesforce Open-source Encoder-decoder Bootstrapping language-image pre-training, designed to bridge vision-language models with large language models for improved visual understanding and generation.
 Truthfulness
-LLaVA1.5-7b/13b 2023 Microsoft Large language and vision assistant models with 7B and 13B parameters, respectively, designed for multimodal tasks by integrating visual information into language models.
+InstructBLIP-7b/13b 2023 Salesforce Open-source Encoder-decoder Visual instruction-tuned versions of BLIP-2, utilizing Vicuna-7B and Vicuna-13B language models, respectively, to enhance vision-language understanding through instruction tuning.
 Truthfulness
-mPLUGOwl2 2023 Zhejiang University Multimodal pre-trained language model designed to handle various vision-language tasks, including image captioning and visual question answering. Truthfulness
-Models Release Year Institution Primary Task Dimensions
-XrayGPT 2023 University of Toronto Specialized model for generating radiology reports from chest X-ray images, aiming to assist in medical image interpretation. Truthfulness
-MiniGPT4 2023 King Abdullah University of Science and Technology A lightweight multimodal model designed to align vision and language models efficiently, facilitating tasks like image captioning and visual question answering.
+LLaVA1.5-7b/13b 2023 Microsoft Open-source Encoder-decoder Large language and vision assistant models with 7B and 13B parameters, respectively, designed for multimodal tasks by integrating visual information into language models.
 Truthfulness
-RadFM 2023 Stanford University Foundation model tailored for radiology, focusing on interpreting medical images and integrating findings with clinical language models. Truthfulness
-Alpaca-LoRA 2023 Stanford University It focuses on achieving good performance in tasks such as question answering and personalized dialogue. Truthfulness
-Robin- medical 2023 Robin Health Fine-tuned for medical applications, including clinical decision support, medical question answering, and health record analysis. Truthfulness
-Flan-T5 2021 Google Research Optimized for tasks like question answering, text summarization, and sentence classification, across a variety of domains. Truthfulness
-BioBERT 2019 Korea University Biomedical language representation learning, enhancing performance on tasks like named entity recognition, relation extraction, and question answering within the biomedical domain.
+mPLUGOwl2 2023 Zhejiang University Open-source Encoder-decoder Multimodal pre-trained language model designed to handle various vision-language tasks, including image captioning and visual question answering.
 Truthfulness
-Falcon Instruct (7B and 40B) 2023 Technology Innovation Institute (TII), UAE. Instruction-tuned language model designed to follow user instructions effectively. Truthfulness, Robustness
-Mistral Instruct (7B) 2023 Mistral AI Instruction-tuned language model designed to follow user instructions effectively. Truthfulness, Robustness
+XrayGPT 2023 University of Toronto Open-source Decoder-only Specialized model for generating radiology reports from chest X-ray images, aiming to assist in medical image interpretation. Truthfulness
+MiniGPT4 2023 King Abdullah University of Science and Technology Open-source Decoder-only A lightweight multimodal model designed to align vision and language models efficiently, facilitating tasks like image captioning and visual question answering.
+Truthfulness
+RadFM 2023 Stanford University Open-source Decoder-only Foundation model tailored for radiology, focusing on interpreting medical images and integrating findings with clinical language models.
+Truthfulness
+Alpaca-LoRA 2023 Stanford University Open-source Decoder-only It focuses on achieving good performance in tasks such as question answering and personalized dialogue. Truthfulness
+Robin- medical 2023 Robin Health Open-source Decoder-only Fine-tuned for medical applications, including clinical decision support, medical question answering, and health record analysis. Truthfulness
+Flan-T5 2021 Google Research Open-source Encoder-decoder Optimized for tasks like question answering, text summarization, and sentence classification, across a variety of domains.
+Truthfulness, Explainability
+BioBERT 2019 Korea University Open-source Encoder-only Biomedical language representation learning, enhancing performance on tasks like named entity recognition, relation extraction, and question answering within the biomedical domain.
+Truthfulness
+Falcon Instruct (7B and 40B) 2023 Technology Innovation Institute (TII), UAE.
+Open-source Decoder-only Instruction-tuned language model designed to follow user instructions effectively.
+Truthfulness, Robustness
+Models Release Year Institution Openness Architecture Primary Task Dimensions
+Mistral Instruct (7B) 2023 Mistral AI Open-source Decoder-only Instruction-tuned language model designed to follow user instructions effectively. Truthfulness, Robustness
 Falcon 2023 Technology Innovation Institute (TII), UAE.
-General-purpose language model optimized for text understanding, generation, question answering, and reasoning tasks.
+Open-source Decoder-only General-purpose language model optimized for text understanding, generation, question answering, and reasoning tasks.
 Focused on efficient deployment for industry-scale applications.
 Truthfulness, Robustness
-LLaVA-Med 2024 Microsoft Large language and vision assistant for biomedicine, trained to handle visual instruction tasks in the biomedical field, aiming for capabilities similar to GPT-4.
+LLaVA-Med 2024 Microsoft Open-source Encoder-decoder Large language and vision assistant for biomedicine, trained to handle visual instruction tasks in the biomedical field, aiming for capabilities similar to GPT-4.
 Truthfulness, Explainability
-GPT-4o-mini 2024 OpenAI Natural language processing (NLP), text generation, and understanding. Explainability
-ASCLEPIUS (13B) 2023 Asclepius AI Medical NLP, clinical text analysis, and healthcare-related tasks. Explainability
-Models Release Year Institution Primary Task Dimensions
-MedViLaM 2023 Xu et al. (2024) Medical vision-language tasks, combining image and text analysis for healthcare. Explainability
-Med-MoE 2023 Jiang et al. (2024) Medical NLP, leveraging Mixture of Experts (MoE) for specialized healthcare tasks. Explainability
-Gemini Pro 2023 Google DeepMind Multi-modal NLP, combining text, image, and other data types for advanced AI tasks Explainability
-AlpaCare (7B) (13B) 2023 Zhang et al. (2023) Healthcare-focused NLP, clinical text analysis, and medical decision support Explainability
-Yi (6B) 2023 01.AI (China) General-purpose NLP, text generation, and fine-tuning for specific applications. Explainability
-Phi-2 (2.7B) 2023 Microsoft Lightweight NLP, text generation, and fine-tuning for specific tasks. Explainability
-SOLAR (10.7B) 2023 Upstage AI General-purpose NLP, text generation, and fine-tuning for specific domains. Explainability
-InternLM2 (7B) 2023 Shanghai AI Laboratory (China) General-purpose NLP, text generation, and fine-tuning for specific applications. Explainability
-Llama3-( 8B and 70B) 2024 Meta General-purpose NLP, text generation, and fine-tuning for specific applications. Privacy
-CodeLlama-( 7B, 13B, and 34B) 2023 Meta Code generation, code completion, and programming assistance. Privacy
-Mixtral-8x7B and 8x22B 2023 Mistral AI General-purpose NLP, text generation, and fine-tuning for specific domains. Privacy
-Qwen-(7B, 14B, 32B, 72B)-Chat 2023 Alibaba Chat-oriented NLP, conversational AI, and text generation. Privacy
-GLM-4 2024 Tsinghua University Advanced NLP, text generation, and multi-modal tasks. Privacy
-Table 2: Detailed Comparison of GPT Models Evaluated for Trust in Healthcare LLMs, Including Model Name, Release Year, Institution, Primary Tasks (e.g., Medical Question-Answering, Clinical Decision Support, Biomedical Text Summarization, Medical Report Generation), and Key Trustworthiness Dimensions (Truthfulness, Privacy, Safety, Robustness, Fairness and Bias, Explainability).
+Claude-3 2024 Anthropic Closed-source Decoder-only General-purpose LLM (QA, dialogue, reasoning, summarization) Explainability
+GPT-4o-mini 2024 OpenAI Closed-source Decoder-only Natural language processing (NLP), text generation, and understanding. Explainability
+ASCLEPIUS (13B) 2023 Asclepius AI Open-source Decoder-only Medical NLP, clinical text analysis, and healthcare-related tasks. Explainability
+MedViLaM 2023 Cite Open-source Encoder-decoder Medical vision-language tasks, combining image and text analysis for healthcare. Explainability
+Med-MoE 2023 Cite Open-source Decoder-only Medical NLP, leveraging Mixture of Experts (MoE) for specialized healthcare tasks. Explainability
+Gemini Pro 2023 Google DeepMind Closed-source Decoder-only Multi-modal NLP, combining text, image, and other data types for advanced AI tasks Explainability
+Gemini-1.5 2024 Google DeepMind Closed-source Decoder-only Multimodal reasoning, long-context understanding, QA, generation Explainability
+AlpaCare (7B) (13B) 2023 Cite Open-source Decoder-only Healthcare-focused NLP, clinical text analysis, and medical decision support Explainability
+Yi (6B) 2023 01.AI (China) Open-source Decoder-only General-purpose NLP, text generation, and fine-tuning for specific applications. Explainability
+Models Release Year Institution Openness Architecture Primary Task Dimensions
+Phi-2 (2.7B) 2023 Microsoft Open-source Decoder-only Lightweight NLP, text generation, and fine-tuning for specific tasks. Explainability
+SOLAR (10.7B) 2023 Upstage AI Open-source Decoder-only General-purpose NLP, text generation, and fine-tuning for specific domains. Explainability
+InternLM2 (7B) 2023 Shanghai AI Laboratory (China) Open-source Decoder-only General-purpose NLP, text generation, and fine-tuning for specific applications. Explainability
+Llama3-( 8B and 70B) 2024 Meta Open-source Decoder-only General-purpose NLP, text generation, and fine-tuning for specific applications. Privacy, Explainability
+CodeLlama-( 7B, 13B, and 34B) 2023 Meta Open-source Decoder-only Code generation, code completion, and programming assistance. Privacy
+Mixtral-8x7B and 8x22B 2023 Mistral AI Open-source Decoder-only General-purpose NLP, text generation, and fine-tuning for specific domains. Privacy
+Qwen-(7B, 14B, 32B, 72B)-Chat 2023 Alibaba Open-source Decoder-only Chat-oriented NLP, conversational AI, and text generation. Privacy
+GLM-4 2024 Tsinghua University Open-source Encoder-decoder Advanced NLP, text generation, and multi-modal tasks. Privacy
+Table 2: Detailed Comparison of GPT Models Evaluated for Trust in Healthcare LLMs, Including Model Name, Release Year, Institution, Openness, Architecture, Primary Tasks (e.g., Medical Question-Answering, Clinical Decision Support, Biomedical Text Summarization, Medical Report Generation), and Key Trustworthiness Dimensions (Truthfulness, Privacy, Safety, Robustness, Fairness and Bias, Explainability).